### v1

# General Chen-Ricci inequalities for Riemannian submersions and Riemannian maps

Abstract
We derive general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. We also derive general forms of the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps between Riemannian manifolds, involving the relations between curvatures of subspaces of source and target spaces. These general forms give new techniques that are easy, elegant, and fruitful to obtain the Chen-Ricci inequalities for such smooth mappings with various space forms. As applications, utilizing these general forms, we explicitly establish Chen-Ricci inequalities when the source manifolds of Riemannian submersions and the target manifolds of Riemannian maps belong to broader classes, such as generalized complex and generalized Sasakian space forms, particularly including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms. Furthermore, we provide Chen-Ricci inequalities for various types of geometrically structured Riemannian submersions and Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant. In particular, imposing suitable circumstances towards validation, our inequalities become the same as various known inequalities established by many authors in [1, 2, 6, 7, 9, 10, 22, 23, 29, 48, 49].
Keywords: Riemannian submersions; Riemannian maps; Ricci curvatures, Chen-Ricci and improved Chen-Ricci inequalities; space forms; complex and contact manifolds.
MSC Classification: 53B20, 53B35, 53C15, 53D15
1 Introduction
In submanifold theory, one of the most fundamental problems is to establish simple and sharp relationships between the intrinsic and extrinsic curvature invariants of a Riemannian submanifold [14]. Toward this, a fundamental relation was derived between the Ricci curvature (intrinsic invariant) and the squared mean curvature (extrinsic invariant) for submanifolds of real space forms [12]. Further, such a relation (known as the Chen-Ricci inequality) has also been established for submanifolds of complex, generalized complex, Sasakian, cosymplectic, Kenmotsu, and generalized Sasakian space forms as listed in Table 1. In addition, in [45], Oprea improved the Chen-Ricci inequality for Lagrangian submanifolds of complex space forms, without determining the equality case, using optimization techniques. In [17], Deng also improved this inequality along with the equality case by using some algebraic inequalities. However, improved Chen-Ricci inequalities for Kählerian slant submanifolds of complex space forms and for $C$-totally real submanifolds of Sasakian space forms cannot be obtained using the approaches of Oprea and Deng. Notably, for such cases, Tripathi established an improved Chen-Ricci inequality with equality in [61]. Since then, many authors have established such inequalities for different types of submanifolds in various kinds of ambient manifolds [34, 30, 33, 35, 39, 40, 41, 46, 61, 67]. For a detailed survey on the Chen-Ricci inequalities, we refer to [15] and the references therein.
Table 1: Chen-Ricci inequalities for submanifolds
Type(s) of Submanifolds Ambient Space Form Ref.
totally real, Lagrangian complex [13]
slant, anti-invariant [37]
slant, invariant, totally real generalized complex [24, 58]
(anti-)invariant, slant Sasakian [26]
$C$-totally real [31, 36, 38]
slant, (anti-)invariant [62]
(anti-)invariant, (semi-)slant [16]
(anti-)invariant Kenmotsu [8]
(anti-)invariant, (semi-)slant cosymplectic [66]
semi-slant, (anti-)invariant [32, 66]
semi-invariant [64]
(anti-)invariant, slant generalized Sasakian [25]
slant [57]
The Chen-Ricci inequalities have also been extensively studied by many authors in the context of geometrically structured Riemannian submersions from such kinds of space forms as listed in Table 2. Motivated by the diverse applications of Riemannian submersions mentioned in [19, 51] and the studies mentioned in Table 2, we introduce a general framework of Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. This framework is suitable to derive Chen-Ricci inequalities for Riemannian submersions when the source manifolds are generalized complex and generalized Sasakian space forms, as well as real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$ space forms, etc. In addition, we can use this general framework to establish Chen-Ricci inequalities for various geometrically structured Riemannian submersions, including invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions having the above-mentioned space forms as the source spaces.
Table 2: Chen-Ricci inequalities for Riemannian submersions
Space Form Type of Submersion Ref.
complex Lagrangian [21]
anti-invariant [23]
slant [22]
hemi-slant [1]
Sasakian anti-invariant [9]
slant [2]
Kenmotsu anti-invariant [48]
slant [6]
cosymplectic anti-invariant [10]
slant [49]
slant generalized Sasakian [7]
The concept of a Riemannian map was introduced by Fischer as a generalization of an isometric immersion and a Riemannian submersion, providing frameworks for quantum models and to compare the geometric properties of two arbitrary Riemannian manifolds [20]. Furthermore, Riemannian maps have been explored in conjunction with complex and contact structures. For detailed studies with applications and examples, see [51] and the references therein. Since Riemannian maps help establish relations between curvature invariants of submanifolds, Lee et al. [29] obtained Chen-Ricci and improved Chen-Ricci inequalities (using Deng’s approach) for Riemannian maps with target spaces as real and complex space forms. Here, we present general forms of the Chen-Ricci and improved Chen-Ricci inequalities (using Tripathi’s approach), along with their equality cases, for Riemannian maps. Notably, these general forms yield the relations between curvatures of subspaces of source and target spaces.
This paper is organized as follows. In Section 2, we provide the definitions of the space forms used in the subsequent sections. Section 3 is dedicated to the Chen-Ricci inequalities for Riemannian submersions. In Subsection 3.1, we provide the necessary preliminaries and foundational concepts required for the subsequent subsections. Subsection 3.2 presents the general forms of Chen-Ricci inequalities for vertical and horizontal distributions. Subsection 3.3 is dedicated to some applications of these general forms. Precisely, we apply these general forms to generalized complex and generalized Sasakian space forms and derive the corresponding Chen-Ricci inequalities, including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms. In addition, we also obtain and validate the inequalities for various types of Riemannian submersions, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions from these spaces. Remarks 3.7 and 3.11 validate the inequalities obtained with the existing ones. Further, Section 4 is dedicated to the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps. In Subsection 4.1, we give some basic notions required for the studies of subsequent subsections. In Subsection 4.2 (resp. Subsection 4.4), we discuss the general form of the Chen-Ricci (resp. improved Chen-Ricci) inequality. Precisely, Theorems 4.1 and 4.7 are devoted to the main results, that is, the general forms of the Chen-Ricci inequalities and the improved Chen-Ricci inequalities for Riemannian maps. Subsections 4.3 and 4.5 are dedicated to some applications of these general forms. Precisely, we focus on the cases when the target spaces of the Riemannian maps are generalized complex and generalized Sasakian space forms. They also cover special cases when the target spaces are real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and almost $C(\alpha)$ space forms. Similarly to the different types of submanifolds, in Theorems 4.5 and 4.10, we also obtain the inequalities for various kinds of Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant to these space forms. Remarks 4.4, 4.6, and 4.11 validate the inequalities obtained with the existing ones. In Section 5, we conclude the paper.
2 Generalized space forms
In this section, we recall the notions of generalized space forms with their special cases.
Generalized complex space forms
Let (M,g,J)({M},g,J) be an even-dimensional Riemannian manifold equipped with an almost complex structure $J$ satisfying $J^{2}=-I$, where $I$ is the identity map. Then (M,g,J)({M},g,J) is called an almost Hermitian manifold if for all ${\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$, the compatibility condition ${g}(J{\cal Y}_{1},J{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})$ holds. Additionally, the manifold is referred to as a generalized complex space form, denoted by ${M}(c_{1},c_{2})$, if there exist smooth functions $c_{1}$ and $c_{2}$ on M{M} such that its curvature tensor $R^{M}$ satisfies [44, 60, 65],
$R^{M}({\cal Y}_{1},{\cal Y}_{2}){\cal Y}_{3}$ $=c_{1}\{g({\cal Y}_{2},{\cal Y}_{3}){\cal Y}_{1}-g({\cal Y}_{1},{\cal Y}_{3}){\cal Y}_{2}\}$ (2.1)
$\quad+c_{2}\{g({\cal Y}_{1},J{\cal Y}_{3})J{\cal Y}_{2}-g({\cal Y}_{2},J{\cal Y}_{3})J{\cal Y}_{1}+2g({\cal Y}_{1},J{\cal Y}_{2})J{\cal Y}_{3}\},$
for all ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. By [63, 65, 60], this includes some special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}$ as listed in Table 3.
Table 3: Special cases of generalized complex space forms
space form $c_{1}$ $c_{2}$
real $c$ 0
complex $\frac{c}{4}$ $\frac{c}{4}$
real Kähler $\frac{c+3\alpha}{4}$ $\frac{c-\alpha}{4}$
Generalized Sasakian space forms
Let $({M},g,\phi,\xi,\eta)$ be an odd-dimensional Riemannian manifold equipped with a $(1,1)$ tensor field $\phi$, a vector field $\xi$, and a $1$-form $\eta$. The structure $({M},g,\phi,\xi,\eta)$ is said to be an almost contact metric manifold if the following conditions are satisfied for all ${\cal Y},{\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$:
$\eta(\xi)=1,~\phi^{2}({\cal Y})=-{\cal Y}+\eta({\cal Y})\xi~{\rm and~}{g}(\phi{\cal Y}_{1},\phi{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})-\eta({\cal Y}_{1})\eta({\cal Y}_{2}).$
Additionally, it holds that $\phi\xi=0$ and $\eta\circ\phi=0$. Furthermore, such a manifold is termed as generalized Sasakian space form (denoted by $M(c_{1},c_{2},c_{3})$) if there exist smooth functions $c_{1}$, $c_{2}$, and $c_{3}$ on $M$ such that the curvature tensor $R^{M}$ of $M$ satisfies [5, 11],
$R^{M}({\cal Y}_{1},{\cal Y}_{2}){\cal Y}_{3}$
$=c_{1}\{g({\cal Y}_{2},{\cal Y}_{3}){\cal Y}_{1}-g({\cal Y}_{1},{\cal Y}_{3}){\cal Y}_{2}\}$
$\quad+c_{2}\{g({\cal Y}_{1},\phi{\cal Y}_{3})\phi{\cal Y}_{2}-g({\cal Y}_{2},\phi{\cal Y}_{3})\phi{\cal Y}_{1}+2g({\cal Y}_{1},\phi{\cal Y}_{2})\phi{\cal Y}_{3}\}$ (2.2)
$\quad+c_{3}\{\eta({\cal Y}_{1})\eta({\cal Y}_{3}){\cal Y}_{2}-\eta({\cal Y}_{2})\eta({\cal Y}_{3}){\cal Y}_{1}+g({\cal Y}_{1},{\cal Y}_{3})\eta({\cal Y}_{2})\xi-g({\cal Y}_{2},{\cal Y}_{3})\eta({\cal Y}_{1})\xi\},$
for any ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. By [5, 11], this also includes several special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}=c_{3}$ as listed in Table 4.
Table 4: Special cases of generalized Sasakian space forms
space form $c_{1}$ $c_{2}=c_{3}$
Sasakian $\frac{c+3}{4}$ $\frac{c-1}{4}$
Kenmotsu $\frac{c-3}{4}$ $\frac{c+1}{4}$
cosymplectic $\frac{c}{4}$ $\frac{c}{4}$
almost $C(\alpha)$ $\frac{c+3\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$
3 General Chen-Ricci inequalities for Riemannian submersions with applications
3.1 Background
In this subsection, we recall some information required for the next subsections.
Let $\left(M_{1}^{m_{1}},g_{1}\right)$ and $\left(M_{2}^{m_{2}},g_{2}\right)$ be two Riemannian manifolds. A surjective smooth map $F:\left(M_{1}^{m_{1}},g_{1}\right)$ $\to\left(M_{2}^{m_{2}},g_{2}\right)$ is called a Riemannian submersion if its differential map (i.e. $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$) is surjective at all $p\in M_{1}$ and $F_{\ast p}$ preserves the length of all horizontal vectors at $p$. Here, the tangent vectors to the fibers (fiber means $\{F^{-1}(q)$ for $q\in M_{2}$}) are referred to as vertical, while the orthogonal vectors to the fibers are termed horizontal. This induces a decomposition of the tangent bundle $TM_{1}$ into a direct sum of two distributions: the vertical distribution $\mathcal{V}=\ker F_{\ast}$ and its orthogonal complement (known as the horizontal distribution) $\mathcal{H}=\left(\ker F_{\ast}\right)^{\perp}$.
O’Neill tensors
For a Riemannian submersion, O’Neill [42] defined two $(1,2)$-type tensor fields T{T} and A{A} that satisfy
${A}_{Y_{1}}Y_{2}=-A_{Y_{2}}Y_{1},$
${T}_{U_{1}}U_{2}={T}_{U_{2}}U_{1},$
$g_{1}\left({T}_{U_{1}}Y_{2},Y_{3}\right)=-g_{1}\left({T}_{U_{1}}Y_{3},Y_{2}\right),$
and
$g_{1}\left({A}_{Y_{1}}Y_{2},Y_{3}\right)=-g_{1}\left({A}_{Y_{1}}Y_{3},Y_{2}\right),$
where $U_{1},U_{2}\in\chi\left(\ker F_{\ast}\right)$ and $Y_{1},Y_{2},Y_{3}\in\chi(\left(\ker F_{\ast}\right)^{\perp})$. Moreover, let $\left\{\vee_{1},\ldots,\vee_{n}\right\}$ be an orthonormal basis of the vertical distribution $(\ker F_{\ast})$. Then the mean curvature vector field $H$ of the fibers of $F$ is defined as [19],
$H=\frac{1}{n}\sum\limits_{i=1}^{n}T_{\vee_{i}}\vee_{i}.$ (3.1)
Relations between Riemannian curvature tensors
Let $R^{M_{1}}$, $R^{M_{2}}$, $R^{\ker F_{\ast}}$, and $R^{\left(\ker F_{\ast}\right)^{\perp}}$ denote the Riemannian curvature tensors corresponding to $M_{1}$, $M_{2}$, $\ker F_{\ast}$, and $\left(\ker F_{\ast}\right)^{\perp}$, respectively. Then we have [42, 19]
$R^{M_{1}}\left(U_{1},U_{2},U_{3},U_{4}\right)$ $=$ $R^{\ker F_{\ast}}\left(U_{1},U_{2},U_{3},U_{4}\right)+g_{1}\left({T}_{U_{1}}U_{4},{T}_{U_{2}}U_{3}\right)$ (3.2)
$-g_{1}\left({T}_{U_{2}}U_{4},{T}_{U_{1}}U_{3}\right),$
$R^{M_{1}}\left(Y_{1},Y_{2},Y_{3},Y_{4}\right)$ $=$ $R^{\left(\ker F_{\ast}\right)^{\perp}}\left(Y_{1},Y_{2},Y_{3},Y_{4}\right)-2g_{1}\left({A}_{Y_{1}}Y_{2},{A}_{Y_{3}}Y_{4}\right)$ (3.3)
$+g_{1}\left({A}_{Y_{2}}Y_{3},{A}_{Y_{1}}Y_{4}\right)-g_{1}\left({A}_{Y_{1}}Y_{3},{A}_{Y_{2}}Y_{4}\right),$
and
$R^{M_{1}}\left(Y_{1},U_{1},Y_{2},U_{2}\right)$ $=$ $g_{1}\left(\left(\nabla_{Y_{1}}^{1}{T}\right)\left(U_{1},U_{2}\right),Y_{2}\right)+g_{1}\left(\left(\nabla_{U_{1}}^{1}{A}\right)\left(Y_{1},Y_{2}\right),U_{2}\right)$ (3.4)
$-g_{1}\left({T}_{U_{1}}Y_{1},{T}_{U_{2}}Y_{2}\right)+g_{1}\left({A}_{Y_{2}}U_{2},{A}_{Y_{1}}U_{1}\right),$
for all $Y_{1},Y_{2},Y_{3},Y_{4}\in\chi\left(\left(\ker F_{\ast}\right)^{\perp}\right)$ and $U_{1},U_{2},U_{3},U_{4}\in\chi\left(\ker F_{\ast}\right)$. Here, $\nabla^{1}$ is the Levi-Civita connection with respect to the metric $g_{1}$.
Some notations
Let $(\ker F_{\ast})={\rm span}\{\vee_{1},\ldots,\vee_{n}\}$ and $\left(\ker F_{\ast}\right)^{\perp}={\rm span}\{h_{1},\ldots,h_{r}\}$. We define some notation as
${T}_{ij}^{t}:=g_{1}({T}_{\vee_{i}}\vee_{j},h_{t}),\quad 1\leq i,j\leq n,\ 1\leq t\leq r,$ (3.5)
$A_{ij}^{\alpha}:=g_{1}({A}_{h_{i}}h_{j},\vee_{\alpha}),\quad 1\leq i,j\leq r,\ 1\leq\alpha\leq n,$ (3.6)
$\delta(N):=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(\left(\nabla^{1}_{h_{i}}{T}\right)_{\vee_{j}}\vee_{j},h_{i}\right),$ (3.7)
$\left\|T^{\cal V}\right\|^{2}:=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}g_{1}\left(T_{\vee_{j}}h_{i},T_{\vee_{j}}h_{i}\right),$ (3.8)
and
$\left\|A^{{\cal H}}\right\|^{2}:=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}g_{1}\left(A_{h_{i}}\vee_{j},A_{h_{i}}\vee_{j}\right).$ (3.9)
Then from [21], we have
$\sum\limits_{t=1}^{r}\sum\limits_{i,j=1}^{n}\left(T_{ij}^{t}\right)^{2}$ $=\frac{1}{2}n^{2}\left\|H\right\|^{2}+\frac{1}{2}\sum\limits_{t=1}^{r}\left(T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right)^{2}$
$+2\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}-2\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}.$ (3.10)
In the sequel, we also use curvature-like notation. Therefore, we fix those as follows.
$2\tau_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}=\sum\limits_{i,j=1}^{r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right),\quad 2\tau_{{\cal H}}^{M_{1}}=\sum\limits_{i,j=1}^{r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right),$
$2\tau_{{\cal V}}^{\ker F_{\ast}}=\sum\limits_{i,j=1}^{n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right),\quad 2\tau_{{\cal V}}^{M_{1}}=\sum\limits_{i,j=1}^{n}R^{M_{1}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right),$
${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})=\sum\limits_{j=1}^{r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{1},h_{j},h_{j},h_{1}\right),\quad{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1})=\sum\limits_{j=1}^{r}R^{M_{1}}\left(h_{1},h_{j},h_{j},h_{1}\right),$
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})=\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}\left(\vee_{1},\vee_{j},\vee_{j},\vee_{1}\right),\quad{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})=\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{1},\vee_{j},\vee_{j},\vee_{1}\right).$
3.2 General framework of Chen-Ricci inequalities
In this subsection, we present general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. Toward this, we have three scenarios depending on the horizontal and vertical distributions. We proceed one by one in the following subsubsections.
3.2.1 For vertical distribution
In this subsubsection, we only consider the vertical distribution. In this case, we have the following result.
Theorem 3.1.
Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})-\frac{1}{4}n^{2}\left\|H\right\|^{2}.$ (3.11)
[…]
$T_{11}^{t}=T_{22}^{t}+\cdots+T_{nn}^{t}\quad\text{and}\quad T_{1j}^{t}=0\quad\text{for}\quad j\in\{2,\ldots,n\},\quad t\in\{1,\ldots,r\}.$
Proof.
For any point $p\in M_{1}$, choose an orthonormal basis of $T_{p}M_{1}$ as $\{h_{1},\ldots,h_{r},\vee_{1},\ldots,\vee_{n}\}$, where $(\ker F_{\ast})={\rm span}\{\vee_{1},\ldots,\vee_{n}\}$ and $\left(\ker F_{\ast}\right)^{\perp}={\rm span}\{h_{1},\ldots,h_{r}\}$. Then from (3.2) and (3.1), we obtain
$2\tau_{{\cal V}}^{\ker F_{\ast}}=2\tau_{{\cal V}}^{M_{1}}-n^{2}\left\|H\right\|^{2}+\sum\limits_{i,j=1}^{n}g_{1}\left(T_{\vee_{j}}\vee_{i},T_{\vee_{i}}\vee_{j}\right).$
Employing (3.5) and then substituting (3.1) yield
$2\tau_{{\cal V}}^{\ker F_{\ast}}$ $=2\tau_{{\cal V}}^{M_{1}}-\frac{1}{2}n^{2}\left\|H\right\|^{2}+\frac{1}{2}\sum\limits_{t=1}^{r}\left(T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right)^{2}$
$\quad+2\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}-2\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}.$
Thus,
$2\tau_{{\cal V}}^{\ker F_{\ast}}\geq 2\tau_{{\cal V}}^{M_{1}}-\frac{1}{2}n^{2}\left\|H\right\|^{2}-2\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}.$ (3.12)
Using (3.2) and (3.5), we derive
$2\sum\limits_{2\leq i<j\leq n}R^{M_{1}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})=$ $2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})$
$+2\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left({T}_{ii}^{t}{T}_{jj}^{t}-({T}_{ij}^{t})^{2}\right).$ (3.13)
Substituting (3.2.1) into (3.12), we get
$2\tau_{{\cal V}}^{\ker F_{\ast}}$ $-2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})$
$\geq 2\tau_{{\cal V}}^{M_{1}}-2\sum\limits_{2\leq i<j\leq n}R^{M_{1}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})-\frac{1}{2}n^{2}\|H\|^{2}.$ (3.14)
Since $2\tau_{{\cal V}}^{\ker F_{\ast}}=2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})+2\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}(\vee_{1},\vee_{j},\vee_{j},\vee_{1})$, the equation (3.2.1) implies (3.11). In addition, the equality case is straightforward by (3.12). ∎
3.2.2 For horizontal distribution
In this subsubsection, we only consider the horizontal distribution. In this case, we have the following result.
Theorem 3.2.
Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1}).$ (3.15)
[…]
$A_{1j}^{\alpha}=0\quad\text{for}\quad j\in\{2,\ldots,r\},\quad\alpha\in\{1,\ldots,n\}.$
Proof.
Utilizing (3.3) and (3.6), we obtain
$\sum\limits_{i,j=1}^{r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)=\sum\limits_{i,j=1}^{r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)-3\sum\limits_{\alpha=1}^{n}\sum\limits_{i,j=1}^{r}\left(A_{ij}^{\alpha}\right)^{2}.$
Equivalently,
$2\tau_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}$ $=2\tau_{{\cal H}}^{M_{1}}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{i,j=1}^{r}\left(A_{ij}^{\alpha}\right)^{2}$
$=2\tau_{{\cal H}}^{M_{1}}-6\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-6\sum\limits_{\alpha=1}^{n}\sum\limits_{2\leq i<j\leq r}\left(A_{ij}^{\alpha}\right)^{2}.$ (3.16)
Again, by substituting $Y_{1}=Y_{4}=h_{i}$ and $Y_{2}=Y_{3}=h_{j}$ in (3.3) and then using (3.6), we obtain
$2\sum\limits_{2\leq i<j\leq r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)=$ $2\sum\limits_{2\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
$+6\sum\limits_{\alpha=1}^{n}\sum\limits_{2\leq i<j\leq r}\left(A_{ij}^{\alpha}\right)^{2}.$ (3.17)
Substituting (3.2.2) into (3.2.2), we deduce that
$2\tau_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}$ $=$ $2\tau_{{\cal H}}^{M_{1}}-6\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
$+2\sum\limits_{2\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)-2\sum\limits_{2\leq i<j\leq r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right).$
Therefore, it follows that
$2~{\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})=2~{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1})-6\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{m}\left(A_{1j}^{\alpha}\right)^{2}.$
This finally gives the desired inequality and the equality case. ∎
3.2.3 For vertical and horizontal distributions
In this subsubsection, we consider both the vertical and horizontal distributions. In this case, we have the following result.
Theorem 3.3.
Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
$\geq{\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)+{\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right)$
$+\delta\left(N\right)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.18)
[…]
$T_{11}^{t}=T_{22}^{t}+\cdots+T_{nn}^{t}\quad\text{and}\quad T_{1j}^{t}=0\quad\text{for}\quad j\in\{2,\ldots,n\},\quad t\in\{1,\ldots,r\}.$
Proof.
By [9, Equation (4.11)], we have the scalar curvature $\tau^{M_{1}}$ of $M_{1}$
$\tau^{M_{1}}=$ $\sum\limits_{1\leq i<j\leq n}R^{M_{1}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
$+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right).$ (3.19)
On the other hand, using (3.2), (3.3) and (3.4), one can get
$2\tau^{M_{1}}$ $=$ $2\tau_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}+2\tau_{{\cal V}}^{\ker F_{\ast}}+n^{2}\left\|H\right\|^{2}$
$+3\sum\limits_{i,j=1}^{r}g_{1}\left(A_{h_{i}}h_{j},A_{h_{i}}h_{j}\right)-\sum\limits_{t=1}^{r}\sum\limits_{i,j=1}^{n}\left(T_{ij}^{t}\right)^{2}$
$-2\sum\limits_{j=1}^{n}\sum\limits_{i=1}^{r}g_{1}\left(\left(\nabla_{h_{i}}^{1}T\right)\left(\vee_{j},\vee_{j}\right),h_{i}\right)$
$+2\sum\limits_{j=1}^{n}\sum\limits_{i=1}^{r}\left\{g_{1}\left(T_{\vee_{j}}h_{i},T_{\vee_{j}}h_{i}\right)-g_{1}\left(A_{h_{i}}\vee_{j},A_{h_{i}}\vee_{j}\right)\right\}.$
Utilizing (3.7), (3.8), (3.9), and (3.1), we rewrite the aforementioned equation as
$\tau^{M_{1}}$ $=\sum\limits_{1\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}$
$\quad-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}$
$\quad+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{2\leq i<j\leq r}\left(A_{ij}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}.$ (3.20)
Using (3.2.1) and (3.2.2) in (3.2.3), we obtain
$\tau^{M_{1}}=$ $\sum\limits_{1\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
$-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
$+\sum\limits_{2\leq i<j\leq n}R^{M_{1}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)-\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)$
$+\sum\limits_{2\leq i<j\leq r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)-\sum\limits_{2\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
$-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}+\frac{n^{2}}{4}\left\|H\right\|^{2}.$ (3.21)
Then by (3.2.3) and (3.2.3), we have
${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)$ $+{\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right)$
$={\rm Ric}^{\ker F_{\ast}}_{\cal V}\left(\vee_{1}\right)+{\rm Ric}^{\left(\ker F_{\ast}\right)^{\perp}}_{\cal H}\left(h_{1}\right)-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]$
$-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}+\frac{n^{2}}{4}\left\|H\right\|^{2}.$
This implies the desired inequality and the equality case. ∎
3.3 Applications
For generalized complex space forms
Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized complex space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)$ using the general inequalities obtained in Subsection 3.2. We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing cases where the source spaces are complex space forms. In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
$J{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
Theorem 3.4.
Let $F:\left(M_{1}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized complex space form onto a Riemannian manifold. Then we have
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2}-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.22)
${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq c_{1}\left(r-1\right)+3c_{2}\left\|Ph_{1}\right\|^{2},$ (3.23)
and
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
$\geq c_{1}\left\{nr+n+r-2\right\}+3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}$
$+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.24)
[…]
Proof.
Let $M_{1}$ be a generalized complex space form. Then, using (2.1), we derive
${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)=c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2},$ (3.25)
${\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)=c_{1}\left(r-1\right)+3c_{2}\left\|Ph_{1}\right\|^{2},$ (3.26)
and
$\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)=nrc_{1}+3c_{2}\|P\|^{2}.$ (3.27)
Adding equations (3.25), (3.26), and (3.27), we obtain
${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)$ $+{\rm Ric}_{{\cal H}}^{{M}_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)$
$=c_{1}\left\{nr+n+r-2\right\}+3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}.$ (3.28)
Then by using (3.25) and (3.11), (3.26) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.22), (3.23) and (3.4) respectively. ∎
Remark 3.5.
[…]
From [52, 53, 54, 55, 47, 59], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost Hermitian manifolds onto Riemannian manifolds. These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the complex structure $J$. Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.4.
Theorem 3.6.
Let $F:\left(M_{1}^{m_{1}}\left(c\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a complex space form onto a Riemannian manifold.
  1. $1.$
    If $F$ is an invariant Riemannian submersion (i.e. $J(\ker F_{\ast})=\ker F_{\ast}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.29)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.30)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}$
    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.31)
  2. $2.$
    If $F$ is an anti-invariant Riemannian submersion (i.e. $J(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.32)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.33)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n+\left\|Ph_{1}\right\|^{2}\right\}$
    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.34)
  3. $3.$
    If $F$ is a semi-invariant Riemannian submersion (i.e. $J(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $J(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  4. $4.$
    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}\cos^{2}\theta-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.35)
    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.36)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}$
    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.37)
  5. $5.$
    If $F$ is a semi-slant Riemannian submersion (i.e. $J({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  6. $6.$
    If $F$ is a hemi-slant Riemannian submersion (i.e. $J({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
Remark 3.7.
In the above theorem, we observe that the corresponding inequalities become the same as in [23, Theorems 4 and 5], [22, Theorem 4.1], and [1, Equations (37), (38), and (62)]. Thus, the inequalities obtained using general forms are validated.
For generalized Sasakian space forms
Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized Sasakian space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)$ using the general inequalities obtained in Subsection 3.2. We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing particular cases where the source spaces are Sasakian, Kenmotsu, cosymplectic, and generalized Sasakian space forms. In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
$\phi{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
Theorem 3.8.
Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold. Then we have
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)$ $+\frac{1}{4}n^{2}\left\|H\right\|^{2}-3c_{2}\left\|Q\vee_{1}\right\|^{2}$
$\geq\left\{\begin{array}[]{l}-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 0,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.40)
${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$ $-c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
$\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.43)
and
${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
$+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}$
$-c_{1}\left\{nr+n+r-2\right\}-3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}$
$\geq\left\{\begin{array}[]{l}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.46)
[…]
Proof.
Let $M_{1}$ be a generalized Sasakian space form. Then, using (2), we derive
${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)=\left\{\begin{array}[]{l}c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2}-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2},~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.49)
${\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)=\left\{\begin{array}[]{l}c_{1}\left(r-1\right)+c_{2}\left\|Ph_{1}\right\|^{2},~{\rm if}~\xi~\text{is vertical};\\ c_{1}\left(r-1\right)+3c_{2}\left\|Ph_{1}\right\|^{2}-c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.52)
and
$\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)=\left\{\begin{array}[]{l}nrc_{1}+3c_{2}\|P\|^{2}-c_{3}r,~{\rm if}~\xi~\text{is vertical};\\ nrc_{1}+3c_{2}\|P\|^{2}-c_{3}n,~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.55)
Adding equations (3.49), (3.52), and (3.55), we obtain
${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)+{\rm Ric}_{{\cal H}}^{{M}_{1}}\left(h_{1}\right)$
$+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)-c_{1}\left\{nr+n+r-2\right\}$
$-3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}$
$=\left\{\begin{array}[]{l}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.58)
Then by using (3.49) and (3.11), (3.52) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.8), (3.8) and (3.8) respectively. ∎
Remark 3.9.
[…]
From [28, 3, 18, 4, 56, 27], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost contact metric manifolds onto Riemannian manifolds. These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the contact structure $\phi$. Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.8.
Theorem 3.10.
Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold.
  1. $1.$
    If $F$ is an invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})=\ker F_{\ast}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
    $\geq\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2},~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.61)
    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$ $-c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.64)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
    $\geq\left\{\begin{array}[]{l}3c_{2}\left\{\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\right\}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}-c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.67)
  2. $2.$
    If $F$ is an anti-invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})$- $c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
    $\geq\left\{\begin{array}[]{l}-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 0,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.70)
    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$- $c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.73)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
    $+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}$
    $-c_{1}\left\{nr+n+r-2\right\}-3c_{2}\left\{n-1+\left\|Ph_{1}\right\|^{2}\right\}$
    $\geq\left\{\begin{array}[]{l}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.76)
  3. $3.$
    If $F$ is a semi-invariant Riemannian submersion (i.e. $\phi(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $\phi(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  4. $4.$
    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}\geq$
    $\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\cos^{2}\theta,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.79)
    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$ $-c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$. (3.82)
    and
    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}$
    $+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
    $\geq\left\{\begin{array}[]{ll}3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta\right\}\\ -c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$. (3.87)
  5. $5.$
    If $F$ is a semi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
  6. $6.$
    If $F$ is a hemi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
    1. $(i)$
[…]
    2. $(ii)$
[…]
Remark 3.11.
If we put the appropriate values of $c_{1}$ and $c_{2}=c_{3}$ from Table 4, then we observe that the corresponding inequalities reduce to the inequalities obtained in [9, 10, Theorems 4.1, 4.2, 4.3, 4.4, 4.5, and 4.6], [48, Theorems 4.1, 4.2, 4.3, 4.6, 4.7, and 4.8], [2, Equations (44), (53), (66), (73), and (74)], [49, Theorems 3.4, 3.5, and 3.6], and [7, Theorems 6, 7, 8, 9, and 10]. Thus, the inequalities obtained using general forms are validated.
4 General Chen-Ricci inequalities for Riemannian maps with applications
4.1 Background
In this subsection, we collect some basic information that is needed for the remaining subsections.
Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a smooth map between Riemannian manifolds with rank $r<\min\{m_{1},m_{2}\}$, and let $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$ be its derivative map at $p$. Denoting the kernel (resp. range) space of $F_{\ast}$ by $(\ker F_{\ast p})$ (resp. $({\rm range~}F_{\ast p})$) and its orthogonal complementary space by $(\ker F_{\ast p})^{\perp}$ (resp. $({\rm range~}F_{\ast p})^{\perp}$), we decompose
$T_{p}M_{1}=(\ker F_{\ast p})\oplus(\ker F_{\ast p})^{\perp}~{\rm and~}T_{F(p)}M_{2}=({\rm range~}F_{\ast p})\oplus({\rm range~}F_{\ast p})^{\perp}.$
The map $F$ is called a Riemannian map, if for all $X,Y\in\Gamma(\ker F_{\ast})^{\perp}$, we have [20]
$g_{1}(X,Y)=g_{2}(F_{\ast}X,F_{\ast}Y).$ (4.1)
Second fundamental form
[…]
$\left(\nabla F_{\ast}\right)\left({\cal Z}_{1},{\cal Z}_{2}\right)=\nabla_{{\cal Z}_{1}}^{F}F_{\ast}({\cal Z}_{2})-F_{\ast}\left(\nabla_{{\cal Z}_{1}}^{{{1}}}{\cal Z}_{2}\right)$
for ${\cal Z}_{1},{\cal Z}_{2}\in\Gamma(T{M_{1}})$, where $\nabla^{F}$ is the pull-back connection. In addition, by [53] $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ is completely contained in $\left({\rm range~}F_{\ast}\right)^{\perp}~{\rm for\ all}\ X,Y\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
Gauss equation
The Gauss equation for $F$ is defined as [51, p. 189]
$g_{2}\left(R^{{M_{2}}}\left(F_{\ast}Z_{1},F_{\ast}Z_{2}\right)F_{\ast}Z_{3},F_{\ast}Z_{4}\right)$ $=$ $g_{1}\left(R^{{M_{1}}}\left(Z_{1},Z_{2}\right)Z_{3},Z_{4}\right)$ (4.2)
$+g_{2}\left(\left(\nabla F_{\ast}\right)\left(Z_{1},Z_{3}\right),\left(\nabla F_{\ast}\right)\left(Z_{2},Z_{4}\right)\right)$
$-g_{2}\left(\left(\nabla F_{\ast}\right)\left(Z_{1},Z_{4}\right),\left(\nabla F_{\ast}\right)\left(Z_{2},Z_{3}\right)\right),$
where $Z_{i}\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$. Here, $R^{M_{1}}$ and $R^{M_{2}}$ denote the curvature tensors of $M_{1}$} and $M_{2}$}, respectively.
Some notations
At a point $p\in{M_{1}}$, suppose that $\{h_{i}\}_{i=1}^{r}$ and $\{h_{i}\}_{i=r+1}^{m_{1}}$ are orthonormal bases of the horizontal space (that is, $(\ker F_{\ast})^{\perp}$) and the vertical space (that is, $(\ker F_{\ast})$), respectively. Then the scalar curvatures $\tau^{{\cal H}}$ and $\tau^{{\cal R}}$ on the horizontal and range spaces are given, respectively, by
$\tau^{{\cal H}}=\sum\limits_{1\leq i<j\leq r}g_{1}\left(R^{{M_{1}}}(h_{i},h_{j})h_{j},h_{i}\right),\quad\tau^{{\cal R}}=\sum\limits_{1\leq i<j\leq r}g_{2}\left(R^{{M_{2}}}(F_{\ast}h_{i},F_{\ast}h_{j})F_{\ast}h_{j},F_{\ast}h_{i}\right).$
In addition, for a fixed $j$, the Ricci curvatures of $h_{j}$ and $F_{\ast}h_{j}$, denoted by ${\rm Ric}(h_{j})$ and ${\rm Ric}(F_{\ast}h_{j})$, respectively, are defined by
${\rm Ric}^{\cal H}(h_{j})=\sum_{i=1}^{r}g_{1}\left(R^{{M_{1}}}(h_{i},h_{j})h_{j},h_{i}\right),{\rm Ric}^{\cal R}(F_{\ast}h_{j})=\sum_{i=1}^{r}g_{2}\left(R^{{M_{2}}}(F_{\ast}h_{i},F_{\ast}h_{j})F_{\ast}h_{j},F_{\ast}h_{i}\right).$ (4.3)
Supposing $\{V_{r+1},\dots,V_{m_{2}}\}$ an orthonormal basis of $\left({\rm range~}F_{\ast}\right)^{\perp}$ we set,
$B_{ij}^{{\cal H}^{\alpha}}$ $=$ $g_{2}\left((\nabla F_{\ast})(h_{i},h_{j}),V_{\alpha}\right),\quad 1\leq i,j\leq r,\quad r+1\leq\alpha\leq m_{2},$
${\rm trace\,}B^{{\cal H}}$ $=$ $\sum_{i=1}^{r}(\nabla F_{\ast})\left(h_{i},h_{i}\right),\quad{\rm and}$
$\left\|{\rm trace\,}B^{{\cal H}}\right\|^{2}$ $=$ $g_{2}\left({\rm trace\,}B^{{\cal H}},{\rm trace\,}B^{{\cal H}}\right).$
Then we get [29],
$\left\|B^{{\cal H}}\right\|^{2}$ $=$ $\sum_{i,j=1}^{r}g_{2}\left((\nabla F_{\ast})(h_{i},h_{j}),(\nabla F_{\ast})(h_{i},h_{j})\right)$ (4.4)
$=$ $\frac{1}{2}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}+\frac{1}{2}\sum_{\alpha=r+1}^{m_{2}}\left(B_{11}^{{\cal H}^{\alpha}}-B_{22}^{{\cal H}^{\alpha}}-\cdots-B_{rr}^{{\cal H}^{\alpha}}\right)^{2}$
$+2\sum_{\alpha=r+1}^{m_{2}}\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{\alpha}}\right)^{2}-2\sum_{\alpha=r+1}^{m_{2}}\sum_{2\leq i<j\leq r}\left\{B_{ii}^{{\cal H}^{\alpha}}B_{jj}^{{\cal H}^{\alpha}}-\left(B_{ij}^{{\cal H}^{\alpha}}\right)^{2}\right\}.$
In the sequel, for any ${\cal Z}\in\Gamma(TM_{2})$, we write
$J{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.5)
and
$\phi{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.6)
where $P{\cal Z}\in\Gamma({\rm range~}F_{\ast})$, $Q{\cal Z}\in\Gamma({\rm range~}F_{\ast})^{\perp}$ and $P$ is the endomorphism on $({\rm range~}F_{\ast})$ such that $\|PF_{\ast}h_{1}\|^{2}=\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}=\sum\limits_{j=1}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}$.
4.2 General framework of Chen-Ricci inequalities
In this subsection, we present a general form of the Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
Theorem 4.1.
Let $F:\left(M_{1}^{m_{1}},g_{1}\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian map between Riemannian manifolds with rank $r<m_{2}$. Then for any unit vector $X\in\Gamma(\ker F_{\ast})^{\perp}$, we have
${\rm Ric}^{{\cal H}}(X)\leq{\rm Ric}^{{\cal R}}(F_{\ast}X)+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$ (4.7)
  1. $(i)$
[…]
    $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ $=$ $0,\quad{\rm for\ all\ }Y\in\Gamma(\ker F_{\ast})^{\perp}\ {\rm orthogonal\ to\ }X,$
    $\left(\nabla F_{\ast}\right)\left(X,X\right)$ $=$ $\frac{{\rm trace~}B^{{\cal H}}}{2}.$
  2. $(ii)$
[…]
Proof.
At a point $p\in{M_{1}}$, let $\left\{h_{1},\dots,h_{r}\right\}$, $\{F_{\ast}h_{1},\dots,F_{\ast}h_{r}\}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$, $({\rm range~}~F_{\ast})$, and $({\rm range~}~F_{\ast})^{\perp}$, respectively. Then by putting $Z_{1}=Z_{4}=h_{i}$ and $Z_{2}=Z_{3}=h_{j}$ in (4.2), we obtain
$2\tau^{{\cal H}}=2\tau^{{\cal R}}+\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}-\left\|B^{{\cal H}}\right\|^{2}.$ (4.8)
Employing (4.4) in (4.8), we get
$\tau^{{\cal H}}$ $=$ $\tau^{{\cal R}}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}-\frac{1}{4}\sum_{\alpha=r+1}^{m_{2}}\left(B_{11}^{{\cal H}^{\alpha}}-B_{22}^{{\cal H}^{\alpha}}-\cdots-B_{rr}^{{\cal H}^{\alpha}}\right)^{2}$ (4.9)
$-\sum_{\alpha=r+1}^{m_{2}}\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{\alpha}}\right)^{2}+\sum_{\alpha=r+1}^{m_{2}}\sum_{2\leq i<j\leq r}\left\{B_{ii}^{{\cal H}^{\alpha}}B_{jj}^{{\cal H}^{\alpha}}-\left(B_{ij}^{{\cal H}^{\alpha}}\right)^{2}\right\}.$
By (4.2), we have
$\sum_{\alpha=r+1}^{m_{2}}\sum_{2\leq i<j\leq r}\left\{B_{ii}^{{\cal H}^{\alpha}}B_{jj}^{{\cal H}^{\alpha}}-\left(B_{ij}^{{\cal H}^{\alpha}}\right)^{2}\right\}$
$=\sum_{2\leq i<j\leq r}g_{1}\left(R^{M_{1}}\left(h_{i},h_{j}\right)h_{j},h_{i}\right)-\sum_{2\leq i<j\leq r}g_{2}\left(R^{M_{2}}\left(F_{\ast}h_{i},F_{\ast}h_{j}\right)F_{\ast}h_{j},F_{\ast}h_{i}\right).$ (4.10)
Then substituting (4.2) in (4.9), we get
$\tau^{{\cal H}}$- $\sum_{2\leq i<j\leq r}g_{1}\left(R^{M_{1}}\left(h_{i},h_{j}\right)h_{j},h_{i}\right)$
$=\tau^{{\cal R}}-\sum_{2\leq i<j\leq r}g_{2}\left(R^{M_{2}}\left(F_{\ast}h_{i},F_{\ast}h_{j}\right)F_{\ast}h_{j},F_{\ast}h_{i}\right)+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$
$-\frac{1}{4}\sum_{\alpha=r+1}^{m_{2}}\left(B_{11}^{{\cal H}^{\alpha}}-B_{22}^{{\cal H}^{\alpha}}-\cdots-B_{rr}^{{\cal H}^{\alpha}}\right)^{2}-\sum_{\alpha=r+1}^{m_{2}}\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{\alpha}}\right)^{2}.$
Equivalently,
${\rm Ric}^{{\cal H}}(h_{1})$ $=$ ${\rm Ric}^{{\cal R}}(F_{\ast}h_{1})+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$ (4.11)
$-\frac{1}{4}\sum_{\alpha=r+1}^{m_{2}}\left(B_{11}^{{\cal H}^{\alpha}}-B_{22}^{{\cal H}^{\alpha}}-\cdots-B_{rr}^{{\cal H}^{\alpha}}\right)^{2}-\sum_{\alpha=r+1}^{m_{2}}\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{\alpha}}\right)^{2}.$
Then (4.7) follows as we can choose $h_{1}=X\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$. Also, the equality holds if and only if
$B_{11}^{{\cal H}^{\alpha}}=B_{22}^{{\cal H}^{\alpha}}+\cdots+B_{rr}^{{\cal H}^{\alpha}},~B_{1i}^{{\cal H}^{\alpha}}=0,\quad i\in\{2,\ldots,r\},\ \alpha\in\{r+1,\ldots,m_{2}\}.$ (4.12)
This implies $(i)$.
Further, assume that the equality case of (4.7) holds for all $X\in\Gamma(\ker F_{\ast})^{\perp}$. Then by (4.12), we have
$B_{ij}^{{\cal H}^{\alpha}}=0\quad~{\rm and}\quad 2B_{ii}^{{\cal H}^{\alpha}}=B_{11}^{{\cal H}^{\alpha}}+B_{22}^{{\cal H}^{\alpha}}+\cdots+B_{rr}^{{\cal H}^{\alpha}},\quad 1\leq i\neq j\leq r.$ (4.13)
This implies
$\left(r-2\right)\sum\limits_{i=1}^{r}B_{ii}^{{\cal H}^{\alpha}}=0.$
Thus, either $\sum\limits_{i=1}^{r}B_{ii}^{{\cal H}^{\alpha}}=0$ or $r=2$.
  • If $r=2$, then using (4.13), we get $B_{11}^{{\cal H}^{\alpha}}=B_{22}^{{\cal H}^{\alpha}}$.
  • If $\sum\limits_{i=1}^{r}B_{ii}^{{\cal H}^{\alpha}}=0$, then using (4.13), we get ${\rm trace~}B^{\cal H}=0$.
We also note that the converse case is trivial. Thus, we complete the required proof. ∎
4.3 Applications
In this subsection, we present the Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms. We also obtain Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms. Furthermore, we derive the Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
In the following theorem, we derive the Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms. Toward this, we use the general form established in Theorem 4.1.
Theorem 4.2.
Suppose $F$ satisfies the hypotheses of Theorem 4.1. Then the following hold.
  1. $(i)$
    If $(M_{2}(c_{1},c_{2}),g_{2},J)$ is a generalized complex space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq T_{GC}=\left(r-1\right)c_{1}+3c_{2}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{GC},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{GC}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{\eta}=\left(r-2\right)\eta\left(F_{\ast}X\right)^{2}$.
In addition, the equality cases are identical to those of Theorem 4.1.
Proof.
Using (4.1), (4.3), (2.1), and (4.5) we obtain,
${\rm Ric}^{{\cal R}}(F_{\ast}h_{1})=\sum\limits_{j=2}^{r}c_{1}\{g_{1}(h_{j},h_{j})g_{1}(h_{1},h_{1})-g_{1}(h_{1},h_{j})g_{1}(h_{j},h_{1})\}$
$+\sum\limits_{j=2}^{r}c_{2}\left\{g_{2}\left(F_{\ast}h_{1},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{j},F_{\ast}h_{1}\right)-g_{2}\left(F_{\ast}h_{j},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{1},F_{\ast}h_{1}\right)\right$.
$\left.+2g_{2}\left(F_{\ast}h_{1},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{j},F_{\ast}h_{1}\right)\right\}.$
Equivalently,
${\rm Ric}^{{\cal R}}(F_{\ast}h_{1})=\left(r-1\right)c_{1}+3c_{2}\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}.$ (4.14)
Then $(i)$ follows by substituting (4.14) in (4.7) as we can choose $h_{1}=X\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
In addition, using (4.1), (4.3), (2), and (4.6) we obtain,
${\rm Ric}^{{\cal R}}(F_{\ast}h_{1})=\sum\limits_{j=2}^{r}c_{1}\left\{g_{1}\left(h_{j},h_{j}\right)g_{1}\left(h_{1},h_{1}\right)-g_{1}\left(h_{1},h_{j}\right)g_{1}\left(h_{j},h_{1}\right)\right\}$
$+\sum\limits_{j=2}^{r}c_{2}\left\{g_{2}\left(F_{\ast}h_{1},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{j},F_{\ast}h_{1}\right)-g_{2}\left(F_{\ast}h_{j},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{1},F_{\ast}h_{1}\right)\right$.
$+\left.2g_{2}\left(F_{\ast}h_{1},PF_{\ast}h_{j}\right)g_{2}\left(PF_{\ast}h_{j},F_{\ast}h_{1}\right)\right\}$
$+\sum\limits_{j=2}^{r}c_{3}\left\{\eta\left(F_{\ast}h_{1}\right)\eta\left(F_{\ast}h_{j}\right)g_{2}\left(F_{\ast}h_{j},F_{\ast}h_{1}\right)-\eta\left(F_{\ast}h_{j}\right)\eta\left(F_{\ast}h_{j}\right)g_{2}\left(F_{\ast}h_{1},F_{\ast}h_{1}\right)\right$.
$\left.+g_{2}\left(F_{\ast}h_{1},F_{\ast}h_{j}\right)\eta\left(F_{\ast}h_{j}\right)\eta\left(F_{\ast}h_{1}\right)-g_{2}\left(F_{\ast}h_{j},F_{\ast}h_{j}\right)\eta\left(F_{\ast}h_{1}\right)\eta\left(F_{\ast}h_{1}\right)\right\}.$
Equivalently,
${\rm Ric}^{{\cal R}}(F_{\ast}h_{1})=$
$\left\{\begin{array}[]{ll}\left(r-1\right)c_{1}+3c_{2}\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}-\left(1+\left(r-2\right)\eta\left(F_{\ast}h_{1}\right)^{2}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast});\\ \left(r-1\right)c_{1}+3c_{2}\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp}.\end{array}\right$. (4.17)
Then $(ii)$ follows by substituting (4.3) in (4.7) as we can choose $h_{1}=X\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$. ∎
Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$. However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$. Therefore, for particular values of $c_{i}$s, we have the following inequalities.
Corollary 4.3.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
  1. $(i)$
    If ${M_{2}}(c)$ is a real space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If ${M_{2}}(c)$ is a complex space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  3. $(iii)$
    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  4. $(iv)$
    If ${M_{2}}(c)$ is a Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  5. $(v)$
    If ${M_{2}}(c)$ is a Kenmotsu space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  6. $(vi)$
    If ${M_{2}}(c)$ is a cosymplectic space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  7. $(vii)$
    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
In addition, the equality cases are identical to those of Theorem 4.2.
Remark 4.4.
We observe that the case $(i)$ in the previous corollary is the same as the case [29,  Theorem  3.1].
Similar to [51, Chapter 6], we can extend the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant, Riemannian maps to a complex (resp. generalized Sasakian) space form. Since these notions impose the same conditions on the distribution $({\rm range~}F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with $J$ (resp. $\phi$), we are directly taking into account these conditions to prove the following result. We also note that since the proof is similar to the proof of Theorem 4.2, we skip it.
Theorem 4.5.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
  1. $(1)$
    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}=\left(r-1\right)\frac{c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}+\frac{3c}{4}\cos^{2}\theta.$
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
  2. $(2)$
    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
In addition, the equality cases are similar to those of Theorem 4.1.
Remark 4.6.
We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.2].
4.4 General framework of improved Chen-Ricci inequalities
In this subsection, we present a general form of the improved Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
Theorem 4.7.
Let $F:\left(M_{1}^{m_{1}},g_{1}\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian map between Riemannian manifolds with rank $r<m_{2}$. Then for any unit vector $X\in\Gamma(\ker F_{\ast})^{\perp}$, we have
${\rm Ric}^{{\cal H}}\left(X\right)\leq{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$ (4.18)
The equality holds for all unit $X\in\Gamma(\ker F_{\ast})^{\perp}$ if and only if either
  1. $(i)$
    $B^{{\cal H}}=0$, or
  2. $(ii)$
    $r=2$ and ( F_) ( h_1,h_1) = 3μV_r+1, ( F_) ( h_2,h_2) = μV_r+1, ( F_) ( h_1,h_2) =μV_r+2
for some function $\mu$ with respect to some orthonormal local frame field $\left\{h_{1},h_{2}\right\}$.
Proof.
At a point $p\in{M_{1}}$, let $\left\{h_{i}\right\}_{i=1}^{r}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$ and $({\rm range~}~F_{\ast})^{\perp}$, respectively, and
$B_{ij}^{{\cal H}^{r+k}}=B_{jk}^{{\cal H}^{r+i}}=B_{ki}^{{\cal H}^{r+j}},B_{ij}^{{\cal H}^{\beta}}=0,\quad i,j,k\in\left\{1,\ldots,r\right\},\beta\in\left\{2r+1,\ldots,m_{2}\right\}.$ (4.19)
By (4.2), we get
${\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)={\rm Ric}^{{\cal H}}\left(X\right)+\sum_{\ell=1}^{r}\sum_{i=2}^{r}\left\{\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}-B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}\right\}.$
Since $\sum\limits_{\ell=1}^{r}\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}\geq\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}+\sum\limits_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}$, the aforementioned equation can be written as,
${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}.$ (4.20)
Using (4.19) in (4.20), we derive
${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{\ell\ell}^{{\cal H}^{r+1}}\right)^{2}.$
[…]
4.5 Applications
In this subsection, we present the improved Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms. We also obtain improved Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms. Furthermore, we derive the improved Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
In the following theorem, we derive the improved Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms. Toward this, we use the general form established in Theorem 4.7.
Theorem 4.8.
Suppose $F$ satisfies the hypotheses of Theorem 4.7. Then the following hold.
  1. $(i)$
    If $(M_{2}(c_{1},c_{2}),g_{2},J)$ is a generalized complex space form, then
    ${\rm Ric}^{{\cal H}}\left(X\right)\leq IT_{GC}=\left(r-1\right)c_{1}+3c_{2}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{GC},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{GC}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{\eta}=\left(r-2\right)\eta\left(F_{\ast}X\right)^{2}$.
In addition, the equality cases are identical to those of Theorem 4.7.
Proof.
The proof follows by using (4.14) and (4.3) in (4.18). ∎
Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$. However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$. Therefore, for particular values of $c_{i}$s, we have the following inequalities.
Corollary 4.9.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
  1. $(i)$
    If ${M_{2}}(c)$ is a real space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If ${M_{2}}(c)$ is a complex space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  3. $(iii)$
    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  4. $(iv)$
    If ${M_{2}}(c)$ is a Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  5. $(v)$
    If ${M_{2}}(c)$ is a Kenmotsu space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  6. $(vi)$
    If ${M_{2}}(c)$ is a cosymplectic space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  7. $(vii)$
    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
In addition, the equality cases are identical to those of Theorem 4.7.
Similarly to Theorem 4.5, taking into account the specific conditions imposed by the notions of Riemannian maps, we can also prove the following result. Since the proof is similar to the proof of Theorem 4.8, we skip it.
Theorem 4.10.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
  1. $(1)$
    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}=\left(r-1\right)\frac{c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}+\frac{3c}{4}\cos^{2}\theta.$
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
  2. $(2)$
    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
In addition, the equality cases are similar to those of Theorem 4.7.
Remark 4.11.
We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.9] when we take $J({\rm range~}F_{\ast})=(\rm range~F_{\ast})^{\perp}$.
5 Conclusion
We have obtained general forms of the Chen-Ricci inequalities for Riemannian submersions and Riemannian maps between Riemannian manifolds. These general forms result in new techniques that are easy, elegant, and fruitful in obtaining the Chen-Ricci inequalities with space forms. As applications of these general forms, we have obtained such inequalities with other space forms, such as real, complex, real Kähler, generalized complex, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$, generalized Sasakian space forms. Moreover, we have validated these inequalities with the verification of existing cases. In the future, they may be explored for other remaining space forms.
Acknowledgment
The author, Kiran Meena, sincerely thanks Prof. Bayram Şahin for his comments that improved the first draft.
Declarations
  • Dedication: K. Meena and K. C. Meena dedicate this article to the welcome of their first child (Madhav Meena).
  • Funding: The author, R. Singh, thanks the Human Resource Development Group (HRDG), Council of Scientific and Industrial Research (CSIR), New Delhi, India, for providing a research fellowship (File No: 09/0013(16054)/2022-EMR-I).
  • Conflict of interest/Competing interests: Not applicable.
  • Ethics approval: The submitted work is original and has not been submitted to more than one journal for simultaneous consideration.
  • Consent for publication and participation: Not applicable.
  • Availability of data, codes, and materials: Not available.
  • Authors’ contributions: All authors have equal contributions.

### v2

@@
 Abstract
-We derive general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds.
-We also derive general forms of the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps between Riemannian manifolds, involving the relations between curvatures of subspaces of source and target spaces.
-These general forms give new techniques that are easy, elegant, and fruitful to obtain the Chen-Ricci inequalities for such smooth mappings with various space forms.
-As applications, utilizing these general forms, we explicitly establish Chen-Ricci inequalities when the source manifolds of Riemannian submersions and the target manifolds of Riemannian maps belong to broader classes, such as generalized complex and generalized Sasakian space forms, particularly including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms.
-Furthermore, we provide Chen-Ricci inequalities for various types of geometrically structured Riemannian submersions and Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant.
-In particular, imposing suitable circumstances towards validation, our inequalities become the same as various known inequalities established by many authors in [1, 2, 6, 7, 9, 10, 22, 23, 29, 48, 49].
-Keywords: Riemannian submersions; Riemannian maps; Ricci curvatures, Chen-Ricci and improved Chen-Ricci inequalities; space forms; complex and contact manifolds.
+In this paper, we derive general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds.
+We also derive general forms of the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps between Riemannian manifolds, involving relations between the curvatures of subspaces of the source and target spaces.
+Further, we illustrate equality cases for all these general forms with two examples.
+These general forms yield new, easy, and elegant techniques that are fruitful in obtaining the Chen-Ricci inequalities for such smooth mappings with various structured manifolds.
+As applications, utilizing these general forms, we explicitly establish Chen-Ricci inequalities when the source manifolds of Riemannian submersions and the target manifolds of Riemannian maps belong to broader classes, such as generalized complex and generalized Sasakian space forms, particularly including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and almost $C(\alpha)$ space forms.
+We also validate our approach by imposing appropriate conditions toward various particular existing cases.
+Keywords: Riemannian submersions; Riemannian maps; (improved) Chen-Ricci inequalities; complex and contact manifolds; space forms.
 MSC Classification: 53B20, 53B35, 53C15, 53D15
 1 Introduction
-In submanifold theory, one of the most fundamental problems is to establish simple and sharp relationships between the intrinsic and extrinsic curvature invariants of a Riemannian submanifold [14].
-Toward this, a fundamental relation was derived between the Ricci curvature (intrinsic invariant) and the squared mean curvature (extrinsic invariant) for submanifolds of real space forms [12].
-Further, such a relation (known as the Chen-Ricci inequality) has also been established for submanifolds of complex, generalized complex, Sasakian, cosymplectic, Kenmotsu, and generalized Sasakian space forms as listed in Table 1.
-In addition, in [45], Oprea improved the Chen-Ricci inequality for Lagrangian submanifolds of complex space forms, without determining the equality case, using optimization techniques.
-In [17], Deng also improved this inequality along with the equality case by using some algebraic inequalities.
+In submanifold or immersion theory, one of the most fundamental problems is to establish simple and sharp relationships between the intrinsic and extrinsic curvature invariants of a Riemannian submanifold [11].
+Toward this, various authors have derived Chen-Ricci inequalities providing relations between the Ricci curvature (intrinsic invariant) and the squared mean curvature (extrinsic invariant) for different types of submanifolds in several space forms.
+These inequalities were also improved by Oprea [26] and Deng [13], using optimization techniques and some algebraic inequalities, respectively.
 However, improved Chen-Ricci inequalities for Kählerian slant submanifolds of complex space forms and for $C$-totally real submanifolds of Sasakian space forms cannot be obtained using the approaches of Oprea and Deng.
-Notably, for such cases, Tripathi established an improved Chen-Ricci inequality with equality in [61].
-Since then, many authors have established such inequalities for different types of submanifolds in various kinds of ambient manifolds [34, 30, 33, 35, 39, 40, 41, 46, 61, 67].
-For a detailed survey on the Chen-Ricci inequalities, we refer to [15] and the references therein.
-Table 1: Chen-Ricci inequalities for submanifolds
-Type(s) of Submanifolds Ambient Space Form Ref.
-totally real, Lagrangian complex [13]
-slant, anti-invariant [37]
-slant, invariant, totally real generalized complex [24, 58]
-(anti-)invariant, slant Sasakian [26]
-$C$-totally real [31, 36, 38]
-slant, (anti-)invariant [62]
-(anti-)invariant, (semi-)slant [16]
-(anti-)invariant Kenmotsu [8]
-(anti-)invariant, (semi-)slant cosymplectic [66]
-semi-slant, (anti-)invariant [32, 66]
-semi-invariant [64]
-(anti-)invariant, slant generalized Sasakian [25]
-slant [57]
-The Chen-Ricci inequalities have also been extensively studied by many authors in the context of geometrically structured Riemannian submersions from such kinds of space forms as listed in Table 2.
-Motivated by the diverse applications of Riemannian submersions mentioned in [19, 51] and the studies mentioned in Table 2, we introduce a general framework of Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds.
-This framework is suitable to derive Chen-Ricci inequalities for Riemannian submersions when the source manifolds are generalized complex and generalized Sasakian space forms, as well as real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$ space forms, etc.
-In addition, we can use this general framework to establish Chen-Ricci inequalities for various geometrically structured Riemannian submersions, including invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions having the above-mentioned space forms as the source spaces.
+To avoid such limitations, Tripathi [40] established an approach to improve Chen-Ricci inequalities.
+For a detailed survey on Chen-Ricci inequalities, we refer to [12] and the references therein.
+We know that a Riemannian submersion is the dual concept of isometric immersion and has various applications [15, 32].
+The concept of a Riemannian map encompasses the concepts of isometric immersion and Riemannian submersion, providing frameworks for quantum models and for comparing the geometric properties of two arbitrary Riemannian manifolds [16, 32].
+Since Riemannian submersions and Riemannian maps help to establish relations between curvature invariants of submanifolds of the source and target manifolds, the Chen-Ricci inequalities have also been studied by many authors in the context of these smooth mappings.
+For geometrically structured Riemannian submersions, they are explored from various space forms, as listed in Table 1.
+In addition, for Riemannian maps with real and complex space forms as the target spaces, Lee et al.
+[22] obtained Chen-Ricci and improved Chen-Ricci inequalities (using Deng’s approach).
 Table 1: Chen-Ricci inequalities for Riemannian submersions
@@
 slant generalized Sasakian [7]
-The concept of a Riemannian map was introduced by Fischer as a generalization of an isometric immersion and a Riemannian submersion, providing frameworks for quantum models and to compare the geometric properties of two arbitrary Riemannian manifolds [20].
-Furthermore, Riemannian maps have been explored in conjunction with complex and contact structures.
-For detailed studies with applications and examples, see [51] and the references therein.
-Since Riemannian maps help establish relations between curvature invariants of submanifolds, Lee et al.
-[29] obtained Chen-Ricci and improved Chen-Ricci inequalities (using Deng’s approach) for Riemannian maps with target spaces as real and complex space forms.
-Here, we present general forms of the Chen-Ricci and improved Chen-Ricci inequalities (using Tripathi’s approach), along with their equality cases, for Riemannian maps.
-Notably, these general forms yield the relations between curvatures of subspaces of source and target spaces.
-This paper is organized as follows.
-In Section 2, we provide the definitions of the space forms used in the subsequent sections.
-Section 3 is dedicated to the Chen-Ricci inequalities for Riemannian submersions.
-In Subsection 3.1, we provide the necessary preliminaries and foundational concepts required for the subsequent subsections.
-Subsection 3.2 presents the general forms of Chen-Ricci inequalities for vertical and horizontal distributions.
-Subsection 3.3 is dedicated to some applications of these general forms.
-Precisely, we apply these general forms to generalized complex and generalized Sasakian space forms and derive the corresponding Chen-Ricci inequalities, including real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and $C(\alpha)$ space forms.
-In addition, we also obtain and validate the inequalities for various types of Riemannian submersions, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant submersions from these spaces.
-Remarks 3.7 and 3.11 validate the inequalities obtained with the existing ones.
-Further, Section 4 is dedicated to the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps.
-In Subsection 4.1, we give some basic notions required for the studies of subsequent subsections.
-In Subsection 4.2 (resp.
-Subsection 4.4), we discuss the general form of the Chen-Ricci (resp.
-improved Chen-Ricci) inequality.
-Precisely, Theorems 4.1 and 4.7 are devoted to the main results, that is, the general forms of the Chen-Ricci inequalities and the improved Chen-Ricci inequalities for Riemannian maps.
-Subsections 4.3 and 4.5 are dedicated to some applications of these general forms.
-Precisely, we focus on the cases when the target spaces of the Riemannian maps are generalized complex and generalized Sasakian space forms.
-They also cover special cases when the target spaces are real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, and almost $C(\alpha)$ space forms.
-Similarly to the different types of submanifolds, in Theorems 4.5 and 4.10, we also obtain the inequalities for various kinds of Riemannian maps, such as invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant to these space forms.
-Remarks 4.4, 4.6, and 4.11 validate the inequalities obtained with the existing ones.
-In Section 5, we conclude the paper.
+Motivatingly, in the current paper, we introduce a general framework of Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds (Theorem 3.1).
+We also introduce general frameworks of the Chen-Ricci and improved Chen-Ricci inequalities (using Tripathi’s approach) for Riemannian maps between Riemannian manifolds (Theorems 4.1 and 4.2).
+Notably, these frameworks unify existing studies and yield relations between the curvatures of subspaces of source and target spaces.
+These frameworks are useful for deriving Chen-Ricci inequalities for Riemannian submersions from arbitrary source manifolds and for Riemannian maps to arbitrary target manifolds (for instance, see Theorems 3.3, 4.4 and 4.5).
+In addition, we can use these general frameworks to establish Chen-Ricci inequalities for various geometrically structured Riemannian submersions and Riemannian maps, such as invariant, anti-invariant, Lagrangian, semi-invariant, slant, semi-slant, hemi-slant, etc.
+(see Section 5).
 2 Generalized space forms
@@
 Generalized complex space forms
-Let (M,g,J)({M},g,J) be an even-dimensional Riemannian manifold equipped with an almost complex structure $J$ satisfying $J^{2}=-I$, where $I$ is the identity map.
-Then (M,g,J)({M},g,J) is called an almost Hermitian manifold if for all ${\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$, the compatibility condition ${g}(J{\cal Y}_{1},J{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})$ holds.
-Additionally, the manifold is referred to as a generalized complex space form, denoted by ${M}(c_{1},c_{2})$, if there exist smooth functions $c_{1}$ and $c_{2}$ on M{M} such that its curvature tensor $R^{M}$ satisfies [44, 60, 65],
+Let (M,g,J)({M},g,J) be an even-dimensional almost Hermitian manifold.
+It is referred to as a generalized complex space form, denoted by ${M}(c_{1},c_{2})$, if its curvature tensor $R^{M}$ satisfies [25, 39, 42],
 $R^{M}({\cal Y}_{1},{\cal Y}_{2}){\cal Y}_{3}$ $=c_{1}\{g({\cal Y}_{2},{\cal Y}_{3}){\cal Y}_{1}-g({\cal Y}_{1},{\cal Y}_{3}){\cal Y}_{2}\}$ (2.1)
-$\quad+c_{2}\{g({\cal Y}_{1},J{\cal Y}_{3})J{\cal Y}_{2}-g({\cal Y}_{2},J{\cal Y}_{3})J{\cal Y}_{1}+2g({\cal Y}_{1},J{\cal Y}_{2})J{\cal Y}_{3}\},$
-for all ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$.
-By [63, 65, 60], this includes some special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}$ as listed in Table 3.
-Table 3: Special cases of generalized complex space forms
-space form $c_{1}$ $c_{2}$
-real $c$ 0
-complex $\frac{c}{4}$ $\frac{c}{4}$
-real Kähler $\frac{c+3\alpha}{4}$ $\frac{c-\alpha}{4}$
+$\quad+c_{2}\{g({\cal Y}_{1},J{\cal Y}_{3})J{\cal Y}_{2}-g({\cal Y}_{2},J{\cal Y}_{3})J{\cal Y}_{1}+2g({\cal Y}_{1},J{\cal Y}_{2})J{\cal Y}_{3}\}$
+for all ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$. Here, $c_{1}$ and $c_{2}$ are smooth functions on M{M}.
 Generalized Sasakian space forms
-Let $({M},g,\phi,\xi,\eta)$ be an odd-dimensional Riemannian manifold equipped with a $(1,1)$ tensor field $\phi$, a vector field $\xi$, and a $1$-form $\eta$.
-The structure $({M},g,\phi,\xi,\eta)$ is said to be an almost contact metric manifold if the following conditions are satisfied for all ${\cal Y},{\cal Y}_{1},{\cal Y}_{2}\in\Gamma(T{M})$:
-$\eta(\xi)=1,~\phi^{2}({\cal Y})=-{\cal Y}+\eta({\cal Y})\xi~{\rm and~}{g}(\phi{\cal Y}_{1},\phi{\cal Y}_{2})=g({\cal Y}_{1},{\cal Y}_{2})-\eta({\cal Y}_{1})\eta({\cal Y}_{2}).$
-Additionally, it holds that $\phi\xi=0$ and $\eta\circ\phi=0$.
-Furthermore, such a manifold is termed as generalized Sasakian space form (denoted by $M(c_{1},c_{2},c_{3})$) if there exist smooth functions $c_{1}$, $c_{2}$, and $c_{3}$ on $M$ such that the curvature tensor $R^{M}$ of $M$ satisfies [5, 11],
+Let $({M},g,\phi,\xi,\eta)$ be an odd-dimensional almost contact metric manifold.
+It is termed as generalized Sasakian space form (denoted by $M(c_{1},c_{2},c_{3})$) if there exist smooth functions $c_{1}$, $c_{2}$, and $c_{3}$ on $M$ such that the curvature tensor $R^{M}$ of $M$ satisfies [5, 10],
 $R^{M}({\cal Y}_{1},{\cal Y}_{2}){\cal Y}_{3}$
@@
 for any ${\cal Y}_{1},{\cal Y}_{2},{\cal Y}_{3}\in\Gamma(T{M})$.
-By [5, 11], this also includes several special space forms with sectional curvature $c$ for particular values of functions $c_{1}$ and $c_{2}=c_{3}$ as listed in Table 4.
-Table 4: Special cases of generalized Sasakian space forms
-space form $c_{1}$ $c_{2}=c_{3}$
-Sasakian $\frac{c+3}{4}$ $\frac{c-1}{4}$
-Kenmotsu $\frac{c-3}{4}$ $\frac{c+1}{4}$
-cosymplectic $\frac{c}{4}$ $\frac{c}{4}$
-almost $C(\alpha)$ $\frac{c+3\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$
+By [5, 10, 41, 42, 39], these generalized space forms include several special space forms based on the particular values of functions $c_{1}$, $c_{2}$, and $c_{3}$ as follows.
+Table 2: Special cases
+space form $c_{1}$ $c_{2}$ $c_{3}$
+real $c$ 0 -
+complex $\frac{c}{4}$ $\frac{c}{4}$ -
+real Kähler $\frac{c+3\alpha}{4}$ $\frac{c-\alpha}{4}$ -
+Sasakian $\frac{c+3}{4}$ $\frac{c-1}{4}$ $\frac{c-1}{4}$
+Kenmotsu $\frac{c-3}{4}$ $\frac{c+1}{4}$ $\frac{c+1}{4}$
+cosymplectic $\frac{c}{4}$ $\frac{c}{4}$ $\frac{c}{4}$
+almost $C(\alpha)$ $\frac{c+3\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$ $\frac{c-\alpha^{2}}{4}$
 3 General Chen-Ricci inequalities for Riemannian submersions with applications
@@
 Let $\left(M_{1}^{m_{1}},g_{1}\right)$ and $\left(M_{2}^{m_{2}},g_{2}\right)$ be two Riemannian manifolds.
-A surjective smooth map $F:\left(M_{1}^{m_{1}},g_{1}\right)$ $\to\left(M_{2}^{m_{2}},g_{2}\right)$ is called a Riemannian submersion if its differential map (i.e.
-$F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$) is surjective at all $p\in M_{1}$ and $F_{\ast p}$ preserves the length of all horizontal vectors at $p$.
-Here, the tangent vectors to the fibers (fiber means $\{F^{-1}(q)$ for $q\in M_{2}$}) are referred to as vertical, while the orthogonal vectors to the fibers are termed horizontal.
+A smooth surjective map $F:\left(M_{1}^{m_{1}},g_{1}\right)$ $\to\left(M_{2}^{m_{2}},g_{2}\right)$ is called a Riemannian submersion if its differential map (i.e., $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$) is surjective at all $p\in M_{1}$ and $F_{\ast p}$ preserves the length of all horizontal vectors at $p$.
+Here, the tangent vectors to the fibers (fiber means $\{F^{-1}(q)$ for $q\in M_{2}$}) are termed vertical, while the orthogonal vectors to the fibers are termed horizontal.
 This induces a decomposition of the tangent bundle $TM_{1}$ into a direct sum of two distributions: the vertical distribution $\mathcal{V}=\ker F_{\ast}$ and its orthogonal complement (known as the horizontal distribution) $\mathcal{H}=\left(\ker F_{\ast}\right)^{\perp}$.
@@
 ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})=\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}\left(\vee_{1},\vee_{j},\vee_{j},\vee_{1}\right),\quad{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})=\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{1},\vee_{j},\vee_{j},\vee_{1}\right).$
-3.2 General framework of Chen-Ricci inequalities
-In this subsection, we present general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds.
-Toward this, we have three scenarios depending on the horizontal and vertical distributions.
-We proceed one by one in the following subsubsections.
-3.2.1 For vertical distribution
-In this subsubsection, we only consider the vertical distribution. In this case, we have the following result.
+3.2 General Chen-Ricci inequalities
+Here, we present general forms of the Chen-Ricci inequalities for Riemannian submersions between Riemannian manifolds. Toward this, we have the following result.
 Theorem 3.1.
 Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
-${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})-\frac{1}{4}n^{2}\left\|H\right\|^{2}.$ (3.11)
+${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq{\rm Ric}_{{\cal V}}^{M_{1}}(\vee_{1})-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.11)
+${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1}),$ (3.12)
+and
+${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
+$\geq{\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)+{\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right)$
+$+\delta\left(N\right)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.13)
+Further, equality cases are as follows.
+  1. $(1)$
 […]
     $T_{11}^{t}=T_{22}^{t}+\cdots+T_{nn}^{t}\quad\text{and}\quad T_{1j}^{t}=0\quad\text{for}\quad j\in\{2,\ldots,n\},\quad t\in\{1,\ldots,r\}.$
+[…]
+    1. $(a)$
+      totally geodesic when $n>2$;
+    2. $(b)$
+      totally umbilical when $n=2$.
+  2. $(2)$
+[…]
+    $A_{1j}^{\alpha}=0\quad\text{for}\quad j\in\{2,\ldots,r\},\quad\alpha\in\{1,\ldots,n\}.$
+[…]
 Proof.
@@
 $+2\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left({T}_{ii}^{t}{T}_{jj}^{t}-({T}_{ij}^{t})^{2}\right).$ (3.15)
-Substituting (3.2.1) into (3.12), we get
+Substituting (3.2) into (3.14), we get
 $2\tau_{{\cal V}}^{\ker F_{\ast}}$ $-2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})$
 $\geq 2\tau_{{\cal V}}^{M_{1}}-2\sum\limits_{2\leq i<j\leq n}R^{M_{1}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})-\frac{1}{2}n^{2}\|H\|^{2}.$ (3.16)
-Since $2\tau_{{\cal V}}^{\ker F_{\ast}}=2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})+2\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}(\vee_{1},\vee_{j},\vee_{j},\vee_{1})$, the equation (3.2.1) implies (3.11).
-In addition, the equality case is straightforward by (3.12).
-∎
-3.2.2 For horizontal distribution
-In this subsubsection, we only consider the horizontal distribution. In this case, we have the following result.
-Theorem 3.2.
-Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
-${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1}).$ (3.15)
-[…]
-$A_{1j}^{\alpha}=0\quad\text{for}\quad j\in\{2,\ldots,r\},\quad\alpha\in\{1,\ldots,n\}.$
-Proof.
-Utilizing (3.3) and (3.6), we obtain
+Since $2\tau_{{\cal V}}^{\ker F_{\ast}}=2\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}(\vee_{i},\vee_{j},\vee_{j},\vee_{i})+2\sum\limits_{j=1}^{n}R^{\ker F_{\ast}}(\vee_{1},\vee_{j},\vee_{j},\vee_{1})$, the equation (3.2) implies (3.11).
+In addition, its equality case is straightforward by (3.14).
+Further, utilizing (3.3) and (3.6), we obtain
 $\sum\limits_{i,j=1}^{r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)=\sum\limits_{i,j=1}^{r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)-3\sum\limits_{\alpha=1}^{n}\sum\limits_{i,j=1}^{r}\left(A_{ij}^{\alpha}\right)^{2}.$
@@
 $+6\sum\limits_{\alpha=1}^{n}\sum\limits_{2\leq i<j\leq r}\left(A_{ij}^{\alpha}\right)^{2}.$ (3.18)
-Substituting (3.2.2) into (3.2.2), we deduce that
+Substituting (3.2) into (3.2), we deduce that
 $2\tau_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}$ $=$ $2\tau_{{\cal H}}^{M_{1}}-6\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
@@
 $2~{\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})=2~{\rm Ric}_{{\cal H}}^{M_{1}}(h_{1})-6\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{m}\left(A_{1j}^{\alpha}\right)^{2}.$
-This finally gives the desired inequality and the equality case. ∎
-3.2.3 For vertical and horizontal distributions
-In this subsubsection, we consider both the vertical and horizontal distributions. In this case, we have the following result.
-Theorem 3.3.
-Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a Riemannian submersion between two Riemannian manifolds. Then
-${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
-$\geq{\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)+{\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right)$
-$+\delta\left(N\right)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.18)
 […]
-$T_{11}^{t}=T_{22}^{t}+\cdots+T_{nn}^{t}\quad\text{and}\quad T_{1j}^{t}=0\quad\text{for}\quad j\in\{2,\ldots,n\},\quad t\in\{1,\ldots,r\}.$
-Proof.
-By [9, Equation (4.11)], we have the scalar curvature $\tau^{M_{1}}$ of $M_{1}$
+Furthermore, by [8, Equation (4.11)], we have the scalar curvature $\tau^{M_{1}}$ of $M_{1}$
 $\tau^{M_{1}}=$ $\sum\limits_{1\leq i<j\leq n}R^{M_{1}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{M_{1}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
@@
 $\tau^{M_{1}}$ $=\sum\limits_{1\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}$
-$\quad-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}$
+$\quad-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+\sum\limits_{t=1}^{r}\sum\limits_{2\leq i<j\leq n}\left\{T_{ii}^{t}T_{jj}^{t}-\left(T_{ij}^{t}\right)^{2}\right\}$
 $\quad+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{2\leq i<j\leq r}\left(A_{ij}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}.$ (3.20)
-Using (3.2.1) and (3.2.2) in (3.2.3), we obtain
+Using (3.2) and (3.2) in (3.2), we obtain
 $\tau^{M_{1}}=$ $\sum\limits_{1\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)+\sum\limits_{1\leq i<j\leq r}R^{\left(\ker F_{\ast}\right)^{\perp}}\left(h_{i},h_{j},h_{j},h_{i}\right)$
-$-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
+$-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
 $+\sum\limits_{2\leq i<j\leq n}R^{M_{1}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)-\sum\limits_{2\leq i<j\leq n}R^{\ker F_{\ast}}\left(\vee_{i},\vee_{j},\vee_{j},\vee_{i}\right)$
@@
 $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}+\frac{n^{2}}{4}\left\|H\right\|^{2}.$ (3.21)
-Then by (3.2.3) and (3.2.3), we have
+Then by (3.2) and (3.2), we have
 ${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)$ $+{\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(h_{i},\vee_{j},\vee_{j},h_{i}\right)$
-$={\rm Ric}^{\ker F_{\ast}}_{\cal V}\left(\vee_{1}\right)+{\rm Ric}^{\left(\ker F_{\ast}\right)^{\perp}}_{\cal H}\left(h_{1}\right)-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]$
+$={\rm Ric}^{\ker F_{\ast}}_{\cal V}\left(\vee_{1}\right)+{\rm Ric}^{\left(\ker F_{\ast}\right)^{\perp}}_{\cal H}\left(h_{1}\right)-\frac{1}{4}\sum\limits_{t=1}^{r}\left[T_{11}^{t}-T_{22}^{t}-\cdots-T_{nn}^{t}\right]^{2}$
 $-\sum\limits_{t=1}^{r}\sum\limits_{j=2}^{n}\left(T_{1j}^{t}\right)^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}+\frac{n^{2}}{4}\left\|H\right\|^{2}.$
-This implies the desired inequality and the equality case. ∎
+[…]
+3.3 Example
+Here, we provide an example satisfying the equality of Chen-Ricci inequalities derived in Theorem 3.1, as follows.
+Example 3.2.
+Considering $\left(M_{1}=(-\frac{\pi}{2},\frac{\pi}{2})\times_{\cos t}\mathbb{C},~g_{1}=dt^{2}+\cos^{2}t\sum\limits_{i=1}^{3}((dx_{i})^{2}+(dy_{i})^{2})\right)$ and $\left(M_{2}=\mathbb{R}^{4},g_{2}=\cos^{2}t\sum\limits_{j=1}^{4}(dz_{j})^{2}\right)$, we define $F:(M_{1},g_{1})\to(M_{2},g_{2})$ such that
+$F(t,x_{1},y_{1},x_{2},y_{2},x_{3},y_{3})=(x_{2},y_{2},x_{3},y_{3}).$
+Then we get
+$(\ker F_{\ast})=\operatorname{span}\left\{\vee_{1}=\frac{\partial}{\partial t},\vee_{2}=\frac{1}{\cos t}\frac{\partial}{\partial x_{1}},\vee_{3}=\frac{1}{\cos t}\frac{\partial}{\partial y_{1}}\right\},$
+$(\ker F_{\ast})^{\perp}=\operatorname{span}\left\{h_{1}=\frac{1}{\cos t}\frac{\partial}{\partial x_{2}},h_{2}=\frac{1}{\cos t}\frac{\partial}{\partial y_{2}},h_{3}=\frac{1}{\cos t}\frac{\partial}{\partial x_{3}},h_{4}=\frac{1}{\cos t}\frac{\partial}{\partial y_{3}}\right\},$
+and
+$({\rm range}~F_{\ast})=\operatorname{span}\left\{F_{\ast}(h_{j})=\frac{1}{\cos t}\frac{\partial}{\partial z_{j}}\right\}_{j=1}^{4},$
+where $\left\{\frac{\partial}{\partial t},\frac{1}{\cos t}\frac{\partial}{\partial x_{1}},\frac{1}{\cos t}\frac{\partial}{\partial y_{1}},\frac{1}{\cos t}\frac{\partial}{\partial x_{2}},\frac{1}{\cos t}\frac{\partial}{\partial y_{2}},\frac{1}{\cos t}\frac{\partial}{\partial x_{3}},\frac{1}{\cos t}\frac{\partial}{\partial y_{3}}\right\}$ and $\left\{\frac{\partial}{\partial z_{j}}\right\}_{j=1}^{4}$ are bases of $T_{p}M_{1}$ and $T_{F(p)}M_{2}$, respectively.
+For all $i,j\in\{1,2,3,4\}$, we verify that
+$g_{1}(h_{i},h_{j})=g_{2}(F_{\ast}h_{i},F_{\ast}h_{j}).$
+Therefore, $F$ is a Riemannian submersion between Riemannian manifolds.
+We calculate the non-vanishing Christoffel symbols for the metric $g_{1}$, which are $\Gamma^{t}_{i,i}=\sin t\cos t$ and $\Gamma^{i}_{t,i}=-\tan t$ for all $i\in\{x_{j},y_{j}\}_{j=1}^{3}$.
+Using these values, we calculate the following:
+$\nabla^{M_{1}}_{\vee_{1}}\vee_{i}=\nabla^{M_{1}}_{\vee_{2}}\vee_{3}=\nabla^{M_{1}}_{\vee_{3}}\vee_{2}=0~\text{for all $1\leq i\leq 3$},$
+$\nabla^{M_{1}}_{\vee_{2}}\vee_{1}=-\tan t\sec t\frac{\partial}{\partial x_{1}},\nabla^{M_{1}}_{\vee_{3}}\vee_{1}=-\tan t\sec t\frac{\partial}{\partial y_{1}},\nabla^{M_{1}}_{\vee_{2}}\vee_{2}=\nabla^{M_{1}}_{\vee_{3}}\vee_{3}=\tan t\frac{\partial}{\partial t},$
+$\nabla^{M_{1}}_{h_{i}}h_{j}=0~\text{for all $1\leq i\neq j\leq 4$ and ~}\nabla^{M_{1}}_{h_{i}}h_{i}=\tan t\frac{\partial}{\partial t}~\text{for all $1\leq i\leq 4$}.$
+Thus, $\mathcal{H}\nabla^{M_{1}}_{\vee_{i}}\vee_{j}=T_{\vee_{i}}\vee_{j}=0$ for all $1\leq i,j\leq 3$. Hence, for all $h_{t}\in\{h_{1},h_{2},h_{3},h_{4}\}$, we have
+$T_{ij}^{t}=g_{1}(T_{\vee_{i}}\vee_{j},h_{t})=0,\quad 1\leq i,j\leq 3~\text{and ~}1\leq t\leq 4.$
+[…]
 3.4 Applications
-For generalized complex space forms
-Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized complex space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)$ using the general inequalities obtained in Subsection 3.2.
-We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing cases where the source spaces are complex space forms.
-In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
-$J{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
-where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
+In this subsection, we obtain the Chen-Ricci inequalities for Riemannian submersions from generalized complex and generalized Sasakian space forms.
+Toward this, we use the general forms established in Theorem 3.1.
+Also, for any ${\cal Y}\in\Gamma(TM_{1})$, we decompose
+$J{\cal Y}{\textrm{~}or~}\phi{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
+where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ with $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
 Theorem 3.3.
-Let $F:\left(M_{1}^{m_{1}}\left(c_{1},c_{2}\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized complex space form onto a Riemannian manifold.
-Then we have
+Suppose $F$ satisfies the hypotheses of Theorem 3.1. Then the following hold.
+  1. $(i)$
+    If $(M_{1}(c_{1},c_{2}),g_{1},J)$ is a generalized complex space form, then
     ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2}-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.22)
@@
     $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.24)
-[…]
-Proof.
-Let $M_{1}$ be a generalized complex space form. Then, using (2.1), we derive
-${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)=c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2},$ (3.25)
-${\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)=c_{1}\left(r-1\right)+3c_{2}\left\|Ph_{1}\right\|^{2},$ (3.26)
-and
-$\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)=nrc_{1}+3c_{2}\|P\|^{2}.$ (3.27)
-Adding equations (3.25), (3.26), and (3.27), we obtain
-${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)$ $+{\rm Ric}_{{\cal H}}^{{M}_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)$
-$=c_{1}\left\{nr+n+r-2\right\}+3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}.$ (3.28)
-Then by using (3.25) and (3.11), (3.26) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.22), (3.23) and (3.4) respectively. ∎
-Remark 3.5.
-[…]
-From [52, 53, 54, 55, 47, 59], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost Hermitian manifolds onto Riemannian manifolds.
-These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the complex structure $J$.
-Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.4.
-Theorem 3.6.
-Let $F:\left(M_{1}^{m_{1}}\left(c\right),g_{1},J\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a complex space form onto a Riemannian manifold.
-  1. $1.$
-    If $F$ is an invariant Riemannian submersion (i.e. $J(\ker F_{\ast})=\ker F_{\ast}$), then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.29)
-    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.30)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
-    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}$
-    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.31)
-  2. $2.$
-    If $F$ is an anti-invariant Riemannian submersion (i.e. $J(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.32)
-    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.33)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
-    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n+\left\|Ph_{1}\right\|^{2}\right\}$
-    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.34)
-  3. $3.$
-    If $F$ is a semi-invariant Riemannian submersion (i.e. $J(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $J(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
-    1. $(i)$
-[…]
   2. $(ii)$
-[…]
-  4. $4.$
-    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})\geq\frac{c}{4}\left(n-1\right)+\frac{3c}{4}\cos^{2}\theta-\frac{1}{4}n^{2}\left\|H\right\|^{2},$ (3.35)
-    ${\rm Ric}_{{\cal H}}^{\left(\ker F_{\ast}\right)^{\perp}}(h_{1})\leq\frac{c}{4}\left(r-1\right)+\frac{3c}{4}\left\|Ph_{1}\right\|^{2},$ (3.36)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)$ $+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
-    $\geq\frac{c}{4}\left\{nr+n+r-2\right\}+\frac{3c}{4}\left\{n\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}$
-    $+\delta(N)-\left\|T^{{\cal V}}\right\|^{2}+\left\|A^{{\cal H}}\right\|^{2}-\frac{n^{2}}{4}\left\|H\right\|^{2}-3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}.$ (3.37)
-  5. $5.$
-    If $F$ is a semi-slant Riemannian submersion (i.e. $J({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
-    1. $(i)$
-[…]
-    2. $(ii)$
-[…]
-  6. $6.$
-    If $F$ is a hemi-slant Riemannian submersion (i.e. $J({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
-    1. $(i)$
-[…]
-    2. $(ii)$
-[…]
-Remark 3.7.
-In the above theorem, we observe that the corresponding inequalities become the same as in [23, Theorems 4 and 5], [22, Theorem 4.1], and [1, Equations (37), (38), and (62)].
-Thus, the inequalities obtained using general forms are validated.
-For generalized Sasakian space forms
-Here, we obtain the Chen-Ricci inequalities for Riemannian submersions whose source spaces are generalized Sasakian space forms $\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)$ using the general inequalities obtained in Subsection 3.2.
-We also provide these inequalities when Riemannian submersions are geometrically structured and validate them with the existing particular cases where the source spaces are Sasakian, Kenmotsu, cosymplectic, and generalized Sasakian space forms.
-In this case, for any ${\cal Y}\in\Gamma(TM_{1})$, we write
-$\phi{\cal Y}={Q}{\cal Y}+{P}{\cal Y},$
-where ${Q}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})$, ${P}{\cal Y}\in\Gamma({\rm ker~}F_{\ast})^{\perp}$ such that $\left\|P\right\|^{2}=\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}\left(g_{1}\left(P\vee_{j},h_{i}\right)\right)^{2}$, $\left\|Ph_{1}\right\|^{2}=\sum\limits_{j=2}^{r}\left(g_{1}\left(Ph_{1},h_{j}\right)\right)^{2}$, and $\left\|Q\vee_{1}\right\|^{2}=\sum\limits_{j=2}^{n}\left(g_{1}\left(Q\vee_{1},\vee_{j}\right)\right)^{2}$.
-Theorem 3.8.
-Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold.
-Then we have
+    If $\left({M_{1}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{1}\right)$ is a generalized Sasakian space form, then
     ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)$ $+\frac{1}{4}n^{2}\left\|H\right\|^{2}-3c_{2}\left\|Q\vee_{1}\right\|^{2}$
@@
 (3.33)
-[…]
-Proof.
-Let $M_{1}$ be a generalized Sasakian space form. Then, using (2), we derive
+In addition, the equality cases are similar to those of Theorem 3.1.
+Proof of $(i)$.
+Using (2.1), we derive
+${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)=c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2},$ (3.34)
+${\rm Ric}_{{\cal H}}^{M_{1}}\left(h_{1}\right)=c_{1}\left(r-1\right)+3c_{2}\left\|Ph_{1}\right\|^{2},$ (3.35)
+and
+$\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)=nrc_{1}+3c_{2}\|P\|^{2}.$ (3.36)
+Adding equations (3.34), (3.35), and (3.36), we obtain
+${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)$ $+{\rm Ric}_{{\cal H}}^{{M}_{1}}\left(h_{1}\right)+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}R^{M_{1}}\left(\vee_{j},h_{i},h_{i},\vee_{j}\right)$
+$=c_{1}\left\{nr+n+r-2\right\}+3c_{2}\left\{\left\|P\right\|^{2}+\left\|Ph_{1}\right\|^{2}+\left\|Q\vee_{1}\right\|^{2}\right\}.$ (3.37)
+Then by using (3.34) and (3.11), (3.35) and (3.12), (3.4) and (3.1), we obtain the desired inequalities (3.22), (3.23) and (i) respectively. ∎
+Proof of $(ii)$.
+Using (2), we derive
 ${\rm Ric}_{{\cal V}}^{M_{1}}\left(\vee_{1}\right)=\left\{\begin{array}[]{l}c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2}-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ c_{1}\left(n-1\right)+3c_{2}\left\|Q\vee_{1}\right\|^{2},~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
@@
 (3.49)
-Then by using (3.49) and (3.11), (3.52) and (3.15), (3.3) and (3.3), we obtain the desired inequalities (3.8), (3.8) and (3.8) respectively. ∎
-Remark 3.9.
-[…]
-From [28, 3, 18, 4, 56, 27], we have the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost contact metric manifolds onto Riemannian manifolds.
-These notions impose different-different conditions on the distribution $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the contact structure $\phi$.
-Using these conditions directly, we can prove the following result along the lines of the proof of Theorem 3.8.
-Theorem 3.10.
-Let $F:\left({M_{1}}^{m_{1}}\left(c_{1},c_{2},c_{3}\right),g_{1},\phi,\xi,\eta\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian submersion from a generalized Sasakian space form onto a Riemannian manifold.
-  1. $1.$
-    If $F$ is an invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})=\ker F_{\ast}$), then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
-    $\geq\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2},~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.61)
-    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$ $-c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
-    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.64)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
-    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
-    $\geq\left\{\begin{array}[]{l}3c_{2}\left\{\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\right\}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{1+\left\|Ph_{1}\right\|^{2}\right\}-c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$.
-(3.67)
-  2. $2.$
-    If $F$ is an anti-invariant Riemannian submersion (i.e. $\phi(\ker F_{\ast})\subseteq(\ker F_{\ast})^{\perp}$), then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})$- $c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}$
-    $\geq\left\{\begin{array}[]{l}-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 0,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.70)
-    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$- $c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
-    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.73)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)$
-    $+\frac{n^{2}}{4}\left\|H\right\|^{2}+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}$
-    $-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}$
-    $-c_{1}\left\{nr+n+r-2\right\}-3c_{2}\left\{n-1+\left\|Ph_{1}\right\|^{2}\right\}$
-    $\geq\left\{\begin{array}[]{l}-c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$.
-(3.76)
-  3. $3.$
-    If $F$ is a semi-invariant Riemannian submersion (i.e. $\phi(\mathcal{D}_{1})=\mathcal{D}_{1}$ and $\phi(\mathcal{D}_{2})\subseteq(\ker F_{\ast})^{\perp}$), and
-    1. $(i)$
-[…]
-    2. $(ii)$
-[…]
-  4. $4.$
-    If $F$ is a slant Riemannian submersion with angle $\theta$, then we have
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}(\vee_{1})-c_{1}\left(n-1\right)+\frac{1}{4}n^{2}\left\|H\right\|^{2}\geq$
-    $\left\{\begin{array}[]{l}3c_{2}\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta-c_{3}\left(1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\cos^{2}\theta,~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.79)
-    ${\rm Ric}_{{\cal H}}^{(\ker F_{\ast})^{\perp}}(h_{1})$ $-c_{1}\left(r-1\right)-3c_{2}\left\|Ph_{1}\right\|^{2}$
-    $\leq\left\{\begin{array}[]{l}0,~{\rm if}~\xi~\text{is vertical};\\ -c_{3}\left(1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right),~{\rm if}~\xi~\text{is horizontal},\end{array}\right$.
-(3.82)
-    and
-    ${\rm Ric}_{{\cal V}}^{\ker F_{\ast}}\left(\vee_{1}\right)+{\rm Ric}_{\cal H}^{(\ker F_{\ast})^{\perp}}\left(h_{1}\right)+\frac{n^{2}}{4}\left\|H\right\|^{2}$
-    $+3\sum\limits_{\alpha=1}^{n}\sum\limits_{j=2}^{r}\left(A_{1j}^{\alpha}\right)^{2}-\delta\left(N\right)+\left\|T^{{\cal V}}\right\|^{2}-\left\|A^{{\cal H}}\right\|^{2}-c_{1}\left\{nr+n+r-2\right\}$
-    $\geq\left\{\begin{array}[]{ll}3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\left(1-\left(\eta\left(\vee_{1}\right)\right)^{2}\right)\cos^{2}\theta\right\}\\ -c_{3}\left\{r+1+\left(n-2\right)\left(\eta\left(\vee_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is vertical};\\ 3c_{2}\left\{\left(n-1\right)\sin^{2}\theta+\left\|Ph_{1}\right\|^{2}+\cos^{2}\theta\right\}\\ -c_{3}\left\{n+1+\left(r-2\right)\left(\eta\left(h_{1}\right)\right)^{2}\right\},&~{\rm if}~\xi~\text{is horizontal}.\end{array}\right$.
-(3.87)
-  5. $5.$
-    If $F$ is a semi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})=\mathcal{D}_{1}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
-    1. $(i)$
-[…]
-    2. $(ii)$
-[…]
-  6. $6.$
-    If $F$ is a hemi-slant Riemannian submersion (i.e. $\phi({\cal D}_{1})\subseteq(\ker F_{\ast})^{\perp}$ and ${\cal D}_{2}$ is slant with angle $\theta$), and
-    1. $(i)$
-[…]
-    2. $(ii)$
-[…]
-Remark 3.11.
-If we put the appropriate values of $c_{1}$ and $c_{2}=c_{3}$ from Table 4, then we observe that the corresponding inequalities reduce to the inequalities obtained in [9, 10, Theorems 4.1, 4.2, 4.3, 4.4, 4.5, and 4.6], [48, Theorems 4.1, 4.2, 4.3, 4.6, 4.7, and 4.8], [2, Equations (44), (53), (66), (73), and (74)], [49, Theorems 3.4, 3.5, and 3.6], and [7, Theorems 6, 7, 8, 9, and 10].
-Thus, the inequalities obtained using general forms are validated.
+Then by using (3.40) and (3.11), (3.43) and (3.12), (3.4) and (3.1), we obtain the desired inequalities (ii), (ii) and (ii) respectively. ∎
 4 General Chen-Ricci inequalities for Riemannian maps with applications
 4.1 Background
-In this subsection, we collect some basic information that is needed for the remaining subsections.
+In this subsection, we recall some information required for the next subsections.
 Let $F:(M_{1}^{m_{1}},g_{1})\to(M_{2}^{m_{2}},g_{2})$ be a smooth map between Riemannian manifolds with rank $r<\min\{m_{1},m_{2}\}$, and let $F_{\ast p}:T_{p}M_{1}\to T_{F(p)}M_{2}$ be its derivative map at $p$.
@@
 […]
-$\left(\nabla F_{\ast}\right)\left({\cal Z}_{1},{\cal Z}_{2}\right)=\nabla_{{\cal Z}_{1}}^{F}F_{\ast}({\cal Z}_{2})-F_{\ast}\left(\nabla_{{\cal Z}_{1}}^{{{1}}}{\cal Z}_{2}\right)$
-for ${\cal Z}_{1},{\cal Z}_{2}\in\Gamma(T{M_{1}})$, where $\nabla^{F}$ is the pull-back connection.
-In addition, by [53] $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ is completely contained in $\left({\rm range~}F_{\ast}\right)^{\perp}~{\rm for\ all}\ X,Y\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
+$\left(\nabla F_{\ast}\right)\left({\cal Z}_{1},{\cal Z}_{2}\right)=\nabla_{F_{\ast}{\cal Z}_{1}}^{2}F_{\ast}({\cal Z}_{2})-F_{\ast}\left(\nabla_{{\cal Z}_{1}}^{{{1}}}{\cal Z}_{2}\right)$ (4.2)
+for ${\cal Z}_{1},{\cal Z}_{2}\in\Gamma(T{M_{1}})$.
+In addition, by [34] $\left(\nabla F_{\ast}\right)\left(X,Y\right)$ is completely contained in $\left({\rm range~}F_{\ast}\right)^{\perp}$ for all $X,Y\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$.
 Gauss equation
-The Gauss equation for $F$ is defined as [51, p. 189]
+The Gauss equation for $F$ is defined as [32, p. 189]
 $g_{2}\left(R^{{M_{2}}}\left(F_{\ast}Z_{1},F_{\ast}Z_{2}\right)F_{\ast}Z_{3},F_{\ast}Z_{4}\right)$ $=$ $g_{1}\left(R^{{M_{1}}}\left(Z_{1},Z_{2}\right)Z_{3},Z_{4}\right)$ (4.3)
@@
 ${\rm Ric}^{\cal H}(h_{j})=\sum_{i=1}^{r}g_{1}\left(R^{{M_{1}}}(h_{i},h_{j})h_{j},h_{i}\right),{\rm Ric}^{\cal R}(F_{\ast}h_{j})=\sum_{i=1}^{r}g_{2}\left(R^{{M_{2}}}(F_{\ast}h_{i},F_{\ast}h_{j})F_{\ast}h_{j},F_{\ast}h_{i}\right).$ (4.4)
-Supposing $\{V_{r+1},\dots,V_{m_{2}}\}$ an orthonormal basis of $\left({\rm range~}F_{\ast}\right)^{\perp}$ we set,
+Supposing $\{V_{r+1},\dots,V_{m_{2}}\}$ an orthonormal basis of $\left({\rm range~}F_{\ast}\right)^{\perp}$ we fix,
 $B_{ij}^{{\cal H}^{\alpha}}$ $=$ $g_{2}\left((\nabla F_{\ast})(h_{i},h_{j}),V_{\alpha}\right),\quad 1\leq i,j\leq r,\quad r+1\leq\alpha\leq m_{2},$
@@
 $+2\sum_{\alpha=r+1}^{m_{2}}\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{\alpha}}\right)^{2}-2\sum_{\alpha=r+1}^{m_{2}}\sum_{2\leq i<j\leq r}\left\{B_{ii}^{{\cal H}^{\alpha}}B_{jj}^{{\cal H}^{\alpha}}-\left(B_{ij}^{{\cal H}^{\alpha}}\right)^{2}\right\}.$
-In the sequel, for any ${\cal Z}\in\Gamma(TM_{2})$, we write
-$J{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.5)
-and
-$\phi{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.6)
-where $P{\cal Z}\in\Gamma({\rm range~}F_{\ast})$, $Q{\cal Z}\in\Gamma({\rm range~}F_{\ast})^{\perp}$ and $P$ is the endomorphism on $({\rm range~}F_{\ast})$ such that $\|PF_{\ast}h_{1}\|^{2}=\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}=\sum\limits_{j=1}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}$.
-4.2 General framework of Chen-Ricci inequalities
-In this subsection, we present a general form of the Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
+4.2 General Chen-Ricci inequality
+Here, we present general form of the Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following result.
 Theorem 4.1.
@@
 […]
+    1. $(a)$
+      totally geodesic when $r>2$;
+    2. $(b)$
+      umbilical when $r=2$.
 Proof.
-At a point $p\in{M_{1}}$, let $\left\{h_{1},\dots,h_{r}\right\}$, $\{F_{\ast}h_{1},\dots,F_{\ast}h_{r}\}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$, $({\rm range~}~F_{\ast})$, and $({\rm range~}~F_{\ast})^{\perp}$, respectively.
-Then by putting $Z_{1}=Z_{4}=h_{i}$ and $Z_{2}=Z_{3}=h_{j}$ in (4.2), we obtain
+Substituting $Z_{1}=Z_{4}=h_{i}$ and $Z_{2}=Z_{3}=h_{j}$ in (4.3), we obtain
 $2\tau^{{\cal H}}=2\tau^{{\cal R}}+\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}-\left\|B^{{\cal H}}\right\|^{2}.$ (4.7)
@@
 We also note that the converse case is trivial. Thus, we complete the required proof. ∎
+4.3 Improved general Chen-Ricci inequality
+In this subsection, we present general form of the improved Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following result.
+Theorem 4.2.
+Let $F:\left(M_{1}^{m_{1}},g_{1}\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian map between Riemannian manifolds with rank $r<m_{2}$.
+Then for any unit vector $X\in\Gamma(\ker F_{\ast})^{\perp}$, we have
+${\rm Ric}^{{\cal H}}\left(X\right)\leq{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$ (4.13)
+The equality holds for all unit $X\in\Gamma(\ker F_{\ast})^{\perp}$ if and only if either
+  1. $(i)$
+    $B^{{\cal H}}=0$, equivalently, $F$ is totally geodesic, or
+  2. $(ii)$
+    $r=2$ and ( F_) ( h_1,h_1) = 3μV_r+1, ( F_) ( h_2,h_2) = μV_r+1, ( F_) ( h_1,h_2) =μV_r+2 for some function $\mu$ with respect to some orthonormal local frame field $\left\{h_{1},h_{2}\right\}$.
+Proof.
+For chosen bases, let
+$B_{ij}^{{\cal H}^{r+k}}=B_{jk}^{{\cal H}^{r+i}}=B_{ki}^{{\cal H}^{r+j}},B_{ij}^{{\cal H}^{\beta}}=0,\quad i,j,k\in\left\{1,\ldots,r\right\},\beta\in\left\{2r+1,\ldots,m_{2}\right\}.$ (4.14)
+By (4.3), we get
+${\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)={\rm Ric}^{{\cal H}}\left(X\right)+\sum_{\ell=1}^{r}\sum_{i=2}^{r}\left\{\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}-B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}\right\}.$
+Since $\sum\limits_{\ell=1}^{r}\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}\geq\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}+\sum\limits_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}$, the aforementioned equation can be written as,
+${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}.$ (4.15)
+Using (4.14) in (4.15), we derive
+${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{\ell\ell}^{{\cal H}^{r+1}}\right)^{2}.$
+[…]
+4.4 Example
+Here, we give an example of a Riemannian map that satisfies the equality of Chen-Ricci and improved Chen-Ricci inequalities derived in Theorems 4.1 and 4.2, as follows.
+Example 4.3.
+Let
+$\left(M_{1}=\left\{(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}:x_{i}\neq 0~\forall~1\leq i\leq 4\right\},g_{1}=e^{2x_{1}}\sum\limits_{i=1}^{2}dx_{i}^{2}+\sum\limits_{i=3}^{4}dx_{i}^{2}\right)$
+and $\left(M_{2}=\left\{(y_{1},y_{2},y_{3},y_{4})\in\mathbb{R}^{4}\right\},g_{2}=e^{2y_{1}}\sum\limits_{j=1}^{2}dx_{j}^{2}+\sum\limits_{j=3}^{4}dx_{j}^{2}\right)$ be two Riemannian manifolds.
+Then by [29], $F:(M_{1},g_{1})\to(M_{2},g_{2})$ defined by
+$F(x_{1},x_{2},x_{3},x_{4})=(x_{1},x_{2},x_{3},0)$
+is a Riemannian map with
+$(\ker F_{\ast})^{\perp}=\operatorname{span}\left\{h_{i}=\frac{\partial}{\partial x_{i}}\right\}_{i=1}^{3},\quad(\ker F_{\ast})=\operatorname{span}\left\{h_{4}=\frac{\partial}{\partial x_{4}}\right\},$
+$({\rm range}~F_{\ast})=\operatorname{span}\left\{F_{\ast}h_{j}=\frac{\partial}{\partial y_{j}}\right\}_{j=1}^{3},\quad({\rm range}~F_{\ast})^{\perp}=\operatorname{span}\left\{\frac{\partial}{\partial y_{4}}\right\},$
+[…]
+[…]
+[…]
 4.5 Applications
-In this subsection, we present the Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms.
-We also obtain Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms.
-Furthermore, we derive the Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
-In the following theorem, we derive the Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms.
-Toward this, we use the general form established in Theorem 4.1.
+In this subsection, we obtain the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms.
+Toward this, we use the general forms established in Theorems 4.1 and 4.2.
+Also, for any ${\cal Z}\in\Gamma(TM_{2})$, we decompose
+$J{\cal Z}~\text{or}~\phi{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.16)
+where $P{\cal Z}\in\Gamma({\rm range~}F_{\ast})$, $Q{\cal Z}\in\Gamma({\rm range~}F_{\ast})^{\perp}$ with
+$\|PF_{\ast}h_{1}\|^{2}=\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}=\sum\limits_{j=1}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}.$
+First, we obtain the Chen-Ricci inequalities as follows.
 Theorem 4.4.
@@
     where $T_{\eta}=\left(r-2\right)\eta\left(F_{\ast}X\right)^{2}$.
-In addition, the equality cases are identical to those of Theorem 4.1.
+The equality conditions are identical to those of Theorem 4.1.
 Proof.
@@
 Then $(ii)$ follows by substituting (4.5) in (4.6) as we can choose $h_{1}=X\in\Gamma\left(\ker F_{\ast}\right)^{\perp}$. ∎
-Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$.
-However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$.
-Therefore, for particular values of $c_{i}$s, we have the following inequalities.
-Corollary 4.3.
-Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
-  1. $(i)$
-    If ${M_{2}}(c)$ is a real space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  2. $(ii)$
-    If ${M_{2}}(c)$ is a complex space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  3. $(iii)$
-    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  4. $(iv)$
-    If ${M_{2}}(c)$ is a Sasakian space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $T_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  5. $(v)$
-    If ${M_{2}}(c)$ is a Kenmotsu space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $T_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  6. $(vi)$
-    If ${M_{2}}(c)$ is a cosymplectic space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $T_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  7. $(vii)$
-    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $T_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-In addition, the equality cases are identical to those of Theorem 4.2.
-Remark 4.4.
-We observe that the case $(i)$ in the previous corollary is the same as the case [29,  Theorem  3.1].
-Similar to [51, Chapter 6], we can extend the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant, Riemannian maps to a complex (resp.
-generalized Sasakian) space form.
-Since these notions impose the same conditions on the distribution $({\rm range~}F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with $J$ (resp.
-$\phi$), we are directly taking into account these conditions to prove the following result.
-We also note that since the proof is similar to the proof of Theorem 4.2, we skip it.
-Theorem 4.5.
-Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
-  1. $(1)$
-    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
-    1. $(i)$
-      $F$ is invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq T_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-    2. $(ii)$
-      $F$ is anti-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}=\left(r-1\right)\frac{c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-    3. $(iii)$
-      $F$ is semi-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-    4. $(iv)$
-      $F$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}+\frac{3c}{4}\cos^{2}\theta.$
-    5. $(v)$
-      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-    6. $(vi)$
-      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-  2. $(2)$
-    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
-    1. $(i)$
-      $F$ is invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    2. $(ii)$
-      $F$ is anti-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    3. $(iii)$
-      $F$ is semi-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    4. $(iv)$
-      $F$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    5. $(v)$
-      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    6. $(vi)$
-      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-In addition, the equality cases are similar to those of Theorem 4.1.
-Remark 4.6.
-We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.2].
-4.4 General framework of improved Chen-Ricci inequalities
-In this subsection, we present a general form of the improved Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
-Theorem 4.7.
-Let $F:\left(M_{1}^{m_{1}},g_{1}\right)\to\left(M_{2}^{m_{2}},g_{2}\right)$ be a Riemannian map between Riemannian manifolds with rank $r<m_{2}$.
-Then for any unit vector $X\in\Gamma(\ker F_{\ast})^{\perp}$, we have
-${\rm Ric}^{{\cal H}}\left(X\right)\leq{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$ (4.18)
-The equality holds for all unit $X\in\Gamma(\ker F_{\ast})^{\perp}$ if and only if either
-  1. $(i)$
-    $B^{{\cal H}}=0$, or
-  2. $(ii)$
-    $r=2$ and ( F_) ( h_1,h_1) = 3μV_r+1, ( F_) ( h_2,h_2) = μV_r+1, ( F_) ( h_1,h_2) =μV_r+2
-for some function $\mu$ with respect to some orthonormal local frame field $\left\{h_{1},h_{2}\right\}$.
-Proof.
-At a point $p\in{M_{1}}$, let $\left\{h_{i}\right\}_{i=1}^{r}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$ and $({\rm range~}~F_{\ast})^{\perp}$, respectively, and
-$B_{ij}^{{\cal H}^{r+k}}=B_{jk}^{{\cal H}^{r+i}}=B_{ki}^{{\cal H}^{r+j}},B_{ij}^{{\cal H}^{\beta}}=0,\quad i,j,k\in\left\{1,\ldots,r\right\},\beta\in\left\{2r+1,\ldots,m_{2}\right\}.$ (4.19)
-By (4.2), we get
-${\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)={\rm Ric}^{{\cal H}}\left(X\right)+\sum_{\ell=1}^{r}\sum_{i=2}^{r}\left\{\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}-B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}\right\}.$
-Since $\sum\limits_{\ell=1}^{r}\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+\ell}}\right)^{2}\geq\sum\limits_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}+\sum\limits_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}$, the aforementioned equation can be written as,
-${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{1\ell}^{{\cal H}^{r+\ell}}\right)^{2}.$ (4.20)
-Using (4.19) in (4.20), we derive
-${\rm Ric}^{{\cal H}}\left(X\right)-{\rm Ric}^{{\cal R}}\left(F_{\ast}X\right)\leq\sum_{\ell=1}^{r}\sum_{i=2}^{r}B_{11}^{{\cal H}^{r+\ell}}B_{ii}^{{\cal H}^{r+\ell}}-\sum_{i=2}^{r}\left(B_{1i}^{{\cal H}^{r+1}}\right)^{2}-\sum_{\ell=2}^{r}\left(B_{\ell\ell}^{{\cal H}^{r+1}}\right)^{2}.$
-[…]
-4.5 Applications
-In this subsection, we present the improved Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms.
-We also obtain improved Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms.
-Furthermore, we derive the improved Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
-In the following theorem, we derive the improved Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms.
-Toward this, we use the general form established in Theorem 4.7.
+Now, we obtain the improved Chen-Ricci inequalities as follows.
 Theorem 4.5.
@@
     where $T_{\eta}=\left(r-2\right)\eta\left(F_{\ast}X\right)^{2}$.
-In addition, the equality cases are identical to those of Theorem 4.7.
+The equality conditions are identical to those of Theorem 4.2.
 Proof.
 The proof follows by using (4.17) and (4.5) in (4.13). ∎
-Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$.
-However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$.
-Therefore, for particular values of $c_{i}$s, we have the following inequalities.
-Corollary 4.9.
-Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
-  1. $(i)$
-    If ${M_{2}}(c)$ is a real space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  2. $(ii)$
-    If ${M_{2}}(c)$ is a complex space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  3. $(iii)$
-    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-  4. $(iv)$
-    If ${M_{2}}(c)$ is a Sasakian space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $IT_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  5. $(v)$
-    If ${M_{2}}(c)$ is a Kenmotsu space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $IT_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  6. $(vi)$
-    If ${M_{2}}(c)$ is a cosymplectic space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $IT_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-  7. $(vii)$
-    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
-    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
-    where $IT_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
-In addition, the equality cases are identical to those of Theorem 4.7.
-Similarly to Theorem 4.5, taking into account the specific conditions imposed by the notions of Riemannian maps, we can also prove the following result.
-Since the proof is similar to the proof of Theorem 4.8, we skip it.
-Theorem 4.10.
-Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
-  1. $(1)$
-    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
-    1. $(i)$
-      $F$ is invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq IT_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-    2. $(ii)$
-      $F$ is anti-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}=\left(r-1\right)\frac{c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
-    3. $(iii)$
-      $F$ is semi-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-    4. $(iv)$
-      $F$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}+\frac{3c}{4}\cos^{2}\theta.$
-    5. $(v)$
-      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-    6. $(vi)$
-      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
-  2. $(2)$
-    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
-    1. $(i)$
-      $F$ is invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    2. $(ii)$
-      $F$ is anti-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    3. $(iii)$
-      $F$ is semi-invariant then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    4. $(iv)$
-      $F$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    5. $(v)$
-      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-    6. $(vi)$
-      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
-      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
-In addition, the equality cases are similar to those of Theorem 4.7.
-Remark 4.11.
-We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.9] when we take $J({\rm range~}F_{\ast})=(\rm range~F_{\ast})^{\perp}$.
-5 Conclusion
-We have obtained general forms of the Chen-Ricci inequalities for Riemannian submersions and Riemannian maps between Riemannian manifolds.
-These general forms result in new techniques that are easy, elegant, and fruitful in obtaining the Chen-Ricci inequalities with space forms.
-As applications of these general forms, we have obtained such inequalities with other space forms, such as real, complex, real Kähler, generalized complex, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$, generalized Sasakian space forms.
-Moreover, we have validated these inequalities with the verification of existing cases.
+5 Conclusion with more applications and validations
+In Theorems 3.1, 4.1 and 4.2, we have obtained general forms of the Chen-Ricci inequalities for Riemannian submersions and Riemannian maps.
+Further, by proving Theorems 3.3, 4.4 and 4.5, we showed that these general forms yield new techniques that are easy, elegant, and fruitful for obtaining Chen-Ricci inequalities with other space forms.
+In view of [25], the item $(i)$ of Theorem 3.3 makes sense for non-constants $c_{1}$ and $c_{2}$ only when $m_{1}=4$.
+Similarly, the item $(i)$ of Theorems 4.4 and 4.5 makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$.
+However, for constants $c_{1}$ and $c_{2}$, the corresponding values of $m_{1}$ or $m_{2}$ could be greater than $4$.
+We note that one can directly obtain Chen-Ricci inequalities with several special space forms by substituting the values of $c_{1}$, $c_{2}$, and $c_{3}$ from Table 2.
+Hence, Theorems 3.3, 4.4 and 4.5 are helpful to explore more particular cases, such as real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, almost $C(\alpha)$ space forms.
+From [33, 34, 35, 36, 27, 38], we have the notions of invariant, anti-invariant, Lagrangian, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost Hermitian manifolds onto Riemannian manifolds.
+In addition, from [21, 3, 14, 4, 37, 20], we have the notions of such structured Riemannian submersions from almost contact metric manifolds onto Riemannian manifolds.
+Similar to [32, Chapter 6], we can also extend the notions of such structured Riemannian maps to a complex or generalized Sasakian space form.
+These notions impose different-different conditions on the decomposition of the distributions $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ and $({\rm range~}F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the complex structure $J$ or contact structure $\phi$.
+Using those conditions directly and putting the appropriate values of $c_{1}$, $c_{2}$ and $c_{3}$, one can also find Chen-Ricci inequalities for such structured Riemannian submersions and Riemannian maps.
+We observe that this approach successfully verifies the following existing inequalities.
+Reference(s) Inequalities
+[1] Equations (37), (38), and (62)
+[2] Equations (44), (53), (66), (73), and (74)
+[7] Theorems 6, 7, 8, 9, and 10
+[8, 9] Theorems 4.1, 4.2, 4.3, 4.4, 4.5, and 4.6
+[18] Theorem 4.1
+[19] Theorems 4 and 5
+[22] Theorems  3.1, 4.2, and 4.3
+[28] Theorems 4.1, 4.2, 4.3, 4.6, 4.7, and 4.8
+[30] Theorems 3.4, 3.5, and 3.6
+Thus, the general forms obtained in this paper are powerful tools towards developing and verifying Chen-Ricci inequalities for Riemannian submersions and Riemannian maps.
 In the future, they may be explored for other remaining space forms.
 Acknowledgment
-The author, Kiran Meena, sincerely thanks Prof. Bayram Şahin for his comments that improved the first draft.
+The second author is grateful to Professor Bayram Şahin for his comments that improved the initial version.
 Declarations
-  • Dedication: K. Meena and K. C. Meena dedicate this article to the welcome of their first child (Madhav Meena).
   • Funding: The author, R.
-Singh, thanks the Human Resource Development Group (HRDG), Council of Scientific and Industrial Research (CSIR), New Delhi, India, for providing a research fellowship (File No: 09/0013(16054)/2022-EMR-I).
+Singh, was supported by the Human Resource Development Group (HRDG), Council of Scientific and Industrial Research (CSIR), New Delhi, India (File No: 09/0013(16054)/2022-EMR-I).
   • Conflict of interest/Competing interests: Not applicable.